### v2

# Beyond English: Evaluating Automated Measurement of Moral Foundations in Non-English Discourse with a Chinese Case Study

Abstract
This study explores computational approaches for measuring moral foundations (MFs) in non-English corpora. Since most resources are developed primarily for English, cross-linguistic applications of moral foundation theory remain limited. Using Chinese as a case study, this paper evaluates the effectiveness of applying English resources to machine translated text, local language lexicons, multilingual language models, and large language models (LLMs) in measuring MFs in non-English texts. The results indicate that machine translation and local lexicon approaches are insufficient for complex moral assessments, frequently resulting in a substantial loss of cultural information. In contrast, multilingual models and LLMs demonstrate reliable cross-language performance with transfer learning, with LLMs excelling in terms of data efficiency. Importantly, this study also underscores the need for human-in-the-loop validation of automated MF assessment, as the most advanced models may overlook cultural nuances in cross-language measurements. The findings highlight the potential of LLMs for cross-language MF measurements and other complex multilingual deductive coding tasks.
Introduction
Moral intuitions have long fascinated social scientists, as they help explain a wide range of cognitive and behavioral phenomena across individuals and groups (Effron and Helgason, 2022). Moral foundation theory (MFT) is among the most prominent psychology frameworks for understanding the origin and development of human morality (Graham et al., 2013). Rooted in moral nativism, MFT argues there are five universal moral foundations—care/harm, fairness/cheating, authority/subversion, loyalty/betrayal, and sanctity/degradation—that transcend languages and cultures and underlie people’s moral judgments and decision-making processes (Graham et al., 2013)  [^1] Each MF includes both virtue and vice dimensions. This paper uses the virtue label to represent the entire foundation.. While some scholars propose other foundations (Haidt, 2012; Atari et al., 2023), these five foundations have received the most empirical validation across domains, languages, and cultures (Iurino and Saucier, 2020).
A growing body of literature employs MFT to investigate online social behaviors. For example, MFT provides a framework for understanding rising political polarization. Individuals who prioritize loyalty, authority, and sanctity foundations are more likely to endorse conservative views and engage in polarized political discourse, while those affiliated with liberal ideologies tend to value all MFs more evenly (Koleva et al., 2012; Haidt and Graham, 2007). MFT also sheds light on online cultural clashes. Atran (2007) identified differing valuations of the sanctity foundation as a key factor in many religious and ideological conflicts. Beyond polarization, MFT has been applied to a range of social issues, including climate change (Markowitz and Shariff, 2012), vaccine hesitancy (Amin et al., 2017), anti-abortion views (Koleva et al., 2012), nationalism (Kertzer et al., 2014), violent collectivism (Nussio, 2023), and terrorism (Tamborini et al., 2020).
Given its broad relevance, measuring MF values in online discourse is essential; yet, automated extraction of MF values from large-scale texts remains challenging, particularly for non-English corpora. Like other latent human values, MFs are often conveyed through abstract narratives that vary across languages. Also, most computational resources for the measurement of MFs are designed for English content (Hoover et al., 2020; Trager et al., 2022). This reliance on English hinders cross-cultural comparative research on MFT and limits theoretical advancement from non-English data (Cheng and Zhang, 2023). Although MFT is a universal theory intended to apply across languages and cultures, the limited availability of non-English resources severely restricts its research scope and further development of the theoretical framework (Graham et al., 2013).
In this work, we investigate various computational approaches for cross-language measurement of MFs with a particular focus on data-efficiency. We use Chinese as an example and find that (1) MF local language lexicons yield suboptimal performance. They are worse than machine translation approaches that utilize established English measurements such as Mformer; (2) multilingual language models can achieve moderate success when trained on annotated English data along with some local language labeled data. However, this strategy is less data-efficient in measuring MFs than in other deductive coding tasks (e.g., Röttger et al., 2022). (3) LLMs outperform other approaches in cross-language MF measurements in both accuracy and data efficiency. Simply fine-tuning and augmenting with English annotated data can achieve strong performance in non-English corpora; (4) nevertheless, this performance is in consistent on different MF values, as LLMs may overlook cultural nuances in cross-language measurements, particularly for culturally distinct values.  [^2] Code is available at GitHub https://github.com/calvinchengyx/cross-lan-mft-measure.
Related Work
The unique link between word usage and the expressed moral values provides a theoretical foundation for automated MF measurement from online texts (Brady, Crockett, and Van Bavel, 2020; Gantman and Van Bavel, 2014, 2016). Scholars have explored different computational approaches, including dictionaries (Graham, Haidt, and Nosek, 2009; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning (Lan and Paraboni, 2022), deep learning language models (Preniqi et al., 2024; Nguyen et al., 2024) and LLMs (Rathje et al., 2024). These computational methods demonstrate great advantages on scalability and labor intensity compared to traditional human annotations.
However, these approaches are primarily developed for English language, and are not directly applicable to non-English texts due to several major concerns: differences in cultural contexts, the lack of annotated datasets, and limitations in domain generalizability. To address the cross-language challenges and bridge knowledge gaps in MF measurements, various computational approaches have been proposed, which can be broadly categorized into two paths: machine translation to English and the development of cross-language measurement tools (Zhuang et al., 2020).
English Centric Machine Translation
Translation is a widely used technique in cross-language MF measurement. For instance, the MFT survey has been translated into over 20 languages for cross-lingual studies (Yilmaz et al., 2016; Nilsson and Erlandsson, 2015). For large scale text analysis, the development of multilingual neural machine translation has significantly improved translation quality compared to earlier statistical methods, enhancing context understanding, ambiguity resolution, and fluency (Stasimioti et al., 2020). This advancement enables a machine-translated approach to cross-language MF measurement by translating target languages into English and applying established English-based methods (Artetxe, Labaka, and Agirre, 2020).
Previous research has proposed various established methods for automatically measuring MFs in English, including dictionaries (Graham, Haidt, and Nosek, 2009; Frimer et al., 2019; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning and deep learning models (Preniqi et al., 2024; Nguyen et al., 2024) trained on annotated English-language social media data (Hoover et al., 2020; Trager et al., 2022).
Moral foundation dictionaries (MFDs)
Word-count methods with crafted English moral lexicons are common. Scholars have developed four English MFDs. The original MFD is an expert-crafted dictionary containing a list of 600 words across five foundation values (Graham, Haidt, and Nosek, 2009). Frimer et al. (2019) then expanded this vocabulary to MFD2 with over 2,000 words by automatically querying similar words with word2vec word embeddings (Mikolov et al., 2013). Similarly, Araque, Gatti, and Kalimeri (2020) extended the original MFD to a MoralStrength dictionary with approximately 1,000 English lemmas based on WordNet synsets (Princeton, 2010). Compared to MFD2, MoralStrength added a round of crowd-sourced ratings on expanded lemmas. Hopp et al. (2021), however, curated a fully crowd-sourced dictionary named eMFD. It is different from previous expert-curated dictionaries for its layperson focus, contextual annotations, probability labeling and large vocabulary with 3,200 words.
Moral word embeddings
Semantic similarity methods using embeddings are another approach. To address the limitations of word-count methods, such as context insensitivity and vocabulary coverage (Nguyen et al., 2024), scholars have introduced semantic similarity methods. For example, Kwak et al. (2021) proposed an embedding framework FrameAxis. It predefines a vector space of micro moral frames with two sets of opposing seed words. Target documents are then converted into vectors using word embedding models, and their MF values are determined by comparing to the micro-frames. This method has been used to extract MF values from various online texts (e.g., Mokhberian et al., 2020; Jing and Ahn, 2021).
Moral language models
Supervised classification models with annotated English-language training data have also been used. With advancements in language models and efforts to create human-labeled MF training datasets (Hoover et al., 2020; Trager et al., 2022), recent studies have demonstrated the potential of fine-tuning language models. For example, Preniqi et al. (2024) fine-tuned a BERT-based classifier MoralBert with large-scale annotated English data and achieved state-of-the-art performance. To address generalizability limitations in out-of-domain datasets (Liscio et al., 2022), Nguyen et al. (2024) proposed another language model Mformer, and reported superior performance compared to other established English methods in evaluations.
Although machine translation offers several advantages in cross-language MF measurement, including interpretability, scalability, efficiency and accessibility, it also faces significant limitations. First, translation quality varies across languages and domains (Ranathunga et al., 2023). In some low-resource languages like Tamil, machine translation often makes errors in translating domain terms, polysemous words, and contains repetitions for semantically similar terms (Ramesh et al., 2021). Second, it often fails to retain non-propositional information, such as emotion nuances. This can lead to emotion loss, toning down, or amplification across languages, introducing bias in subsequent analyses (Troiano, Klinger, and Padó, 2020). Third, machine translation struggles to capture cultural elements, which is a major concern for cross-cultural and comparative research (Haidt, 2012). It often shows limited performance with rare or culture-specific words, idiomatic phrases, and metaphor recognition (Dorothy, Keiko, and Vladimir, 2019). Thus, it remains unclear whether machine translation is a reliable method for measuring cross-language MFs.
Cross-language Measurement Tools
A second path is to develop cross-language tools, where scholars create computational MF resources tailored to local languages. Common cross-language measurements include local language dictionaries, task-specific multilingual language models, and LLMs.
Local language dictionaries
Due to the efficiency at scale and multilingual capabilities, local language dictionaries are widely used to estimate MF values from non-English texts (Hopp et al., 2021). Developing local language MFDs generally involves three steps: (1) translating English MFDs to target languages; (2) adding culturally specific and non-translatable vocabulary; and (3) validating with native speakers and local language corpora. Several extensive non-English MFDs have been developed and validated in Turkish (Alper et al., 2020), Japanese (Matsuo et al., 2019), Portuguese (Carvalho et al., 2020), and Chinese (Cheng and Zhang, 2023). Despite the abovementioned advantages, the dictionary approach still faces the inherent limitations of general bag-of-words methods. In this paper, we use C-MFD2—a Chinese MFD—as a cross-language tool to evaluate a local language dictionary approach. We also test a semantic similarity approach using the FrameAxis architecture and cross-language word embedding models.
Multilingual Language Models
To overcome the limitations of bag-of-words methods, literature has suggested machine learning and deep learning approaches. A primary challenge with these methods is the scarcity of annotated data in local languages for model training (e.g., Ji et al., 2024; Nguyen et al., 2024). Therefore, scholars have adopted transfer learning techniques that leverage English-annotated resources for cross-language classifier development. Two major transfer learning strategies are commonly proposed: (1) machine-translating annotated English-language data into local languages to train monolingual models (Schuster et al., 2019), or (2) using annotated English-language data to train multilingual models (Barriere and Balahur, 2020). Given the strong performance of multilingual language models in deductive coding tasks such as sentiment analysis (e.g., Barriere and Balahur, 2020) and hate speech detection (e.g., Röttger et al., 2022), this paper focuses on the second transfer learning strategy and evaluates multilingual language models for cross-language MF measurement.
Large Language Models (LLMs)
The rise of LLMs provides an alternative for cross-language MF measurement. LLMs show exceptional zero/few-shot learning capability, enabling them to directly label human values out-of-the-box, which is particularly valuable for tasks with limited human annotated data (Ziems et al., 2024). They also have demonstrated strong performance in measuring various human values, including emotions, stance, and political ideology (Ziems et al., 2024; Gilardi, Alizadeh, and Kubli, 2023; Törnberg, 2023). Notably, LLMs sometimes are not as good as specialized fine-tuned language models (Amin, Cambria, and Schuller, 2023; Preniqi et al., 2024), which may be due to the lack of explicit, colloquial definitions of the target human values (Ziems et al., 2024). MFT’s well-established conceptual framework may help address this limitation. Not only are LLMs pre-trained on rich MFT literature (e.g., Abdulhai et al., 2023), but MFT also offers clear guidance for crafting clear and effective prompts. Additionally, LLMs trained on vast multilingual data exhibit promising capabilities to handle cross-language measurements (Ahuja et al., 2023).
Despite the strengths in accessibility, efficiency, multilingualism, and reasoning, there are also some concerns for LLMs’ performance in cross-language MF measurement. First, LLMs exhibit a substantial degree of subordinate multilingualism, displaying proficiency in some languages but not others (Zhang et al., 2023), which has a strong correlation with the proportion of those languages in the pre-training corpus (Li et al., 2024). Second, there are potential language biases, particularly in human-value relevant coding tasks (Kirk et al., 2024). For example, non-English prompts are more likely to generate malicious responses compared to English prompts (Shen et al., 2024). LLMs may also have moral bias across languages, as pre-trained multilingual language models often display distinct moral directions across languages (Hämmerl et al., 2023). Third, different LLMs exhibit varying baseline moral tendencies (Ji et al., 2024). For instance, GPT-3’s MF preferences align more closely with politically conservative individuals when minimal prompt engineering is used in zero-shot learning (Abdulhai et al., 2023). These moral tendencies, however, are highly sensitive to prompting, with different prompting strategies significantly influencing classification outcomes in MF measurements (Abdulhai et al., 2023).
Thus, it is unclear how LLMs performs in cross-language MF measurement tasks. This paper selects a cutting-edge, open-source model—Llama3.1 (Meta, 2024) to evaluate the LLMs approach.
Data
Table 1 shows details of the benchmarking and training datasets used in this work. We used three human annotated datasets—moral foundation vignettes (MFV), Chinese moral scenarios (CCS) and Chinese core values (CCV), to benchmark the performance of different cross-language MF measurement approaches. And we used three English annotated MFT datasets—Reddit Corpus, Twitter Corpus and news corpus—for model training and fine-tuning, aligning with the work of Nguyen et al. (2024).
MFV CCS CCV EN Total
care/harm 27 389 3,030 16,607 20,053
loyalty/betrayal 16 248 1,712 11,772 13,748
authority/subversion 25 331 1,278 13,176 14,810
fairness/cheating 12 259 1,225 16,292 17,788
sanctity/degradation 10 226 247 9,165 9,648
non-moral 0 0 0 28,988 28,988
Total 90 1,453 7,492 71,242 80,277
Table 1: Moral foundation annotated datasets used in this paper. MFV, CCS and 20% of CCV were used for benchmarking; EN and 80% of CCV were used for fine-tuning XLM-T and Llama3.1-8b language models.
Moral Foundation Vignettes (MFV)
MFV is a list of social behaviors constructed by psychologists based on MFT, describing the violations of specific moral values from a third-party perspective (Clifford et al., 2015). It has been widely validated and used to assess measures of moral judgment (e.g., Atari et al., 2023; Curry, Chesters, and Van Lissa, 2019). The original MFV is in English; we manually translate it to Chinese and ensure the cultural context is preserved. MFV is used as a universal stimulus to benchmark the performance of different cross-language MF measurement approaches on an expert-crafted sample.
Chinese Moral Scenarios (CCS)
CCS is list of moral scenarios written by laypeople to describe their intuitive understanding of MF values (Cheng and Zhang, 2023). It is written by 202 native Chinese speakers describing situations that either violate or respect specific MF values, with minimal training in MFT literacy. This reverse-annotation method is commonly used for validating MF measurements (e.g., Cheng and Zhang, 2023; Frimer et al., 2019; Matsuo et al., 2019). It incorporates culturally specific content, reflecting native speakers’ natural and intuitive understanding of MFT in a real-word context.
Chinese Core Values (CCV)
CCV is a human-annotated real-world dataset, including 6,994 sentences collected from four local Chinese news websites. [^3] CCV includes data from the CMOS corpus (Peng et al., 2021), China Cultural and Ethical Website wenming.cn, Youth Patriotism News agzy.youth.cn, and Sohu News news.sohu.com. The dataset is annotated by three native Chinese speakers based on the Chinese core socialist moral value coding scheme, which is highly correlate with the five universal moral foundation values (Liu et al., 2022). The original CCV dataset includes eight labels, [^4] The Chinese core socialist moral values include civility, justice, equality, rule of law, patriotism, dedication, integrity, and friendship. The curated mapping scheme to the five moral foundation values is in the Appendix which are qualitatively evaluated and re-mapped to the universal five foundation values following an expert-curated mapping scheme. We sample 20% of the CCV as the primary benchmarking dataset to evaluate the performance across approaches stratifying on the values. The remaining 80% is used as local-language annotated data to test the data-efficiency fine-tuning language models.
All benchmarking documents are single-class labeled. For multi-class predicted documents, we decided the measurement performance by a lenient evaluation criterion: a prediction is considered correct if one of the predicted values matches the true label.
English annotated data (EN)
We use three English annotated MF datasets for transfer learning, including a Twitter corpus (Hoover et al., 2020), a Reddit corpus (Trager et al., 2022) and a news corpus (Hopp et al., 2021), which are widely used in training English MF classifiers and show reliable performance (Nguyen et al., 2024; Preniqi et al., 2024).
Methods
Machine Translation
We use Google Translate as an example of a machine translation approach due to its accessibility and consistent performance across domains. First, we machine translate benchmarking datasets except MFV to Chinese using the Google Cloud API—Basic Translation service. Then we estimate the MF values from translated documents with established English measurements, including lexicons MFD (Graham, Haidt, and Nosek, 2009), MFD2 (Frimer et al., 2019), eMFD (Hopp et al., 2021) and MoralStrength (Araque, Gatti, and Kalimeri, 2020); word embeddings with FrameAxis (Kwak et al., 2021); and specialized-fine-tuned language models MoralBert (Preniqi et al., 2024) and Mformer (Nguyen et al., 2024).
For MFD, MFD2, and eMFD, we calculate word frequencies using the eMFDscore Python package (Hopp et al., 2021). For MFD and MFD 1.0, each document’s MF value is determined by the most frequent MF class in the respective dictionary. A document is mapped to multiple classes if there is an equal number of class matches. If there are no matching words, no class is assigned to the document. eMFD, however, assigns probabilities to its vocabulary, representing their likelihood of being associated with certain MF classes. We sum the probabilities and label the document with the class that has the highest sum.
We use the moralstrength package (Araque, Gatti, and Kalimeri, 2020) for the MoralStrength dictionary. We first test its performance of the lexicon features alone with bag-of-word methods; then we train a Support Vector Machine (SVM) model and combine its lexicon features. Since English training sets contain many non-moral labels, while the benchmarking dataset contains only moral labels, we train two SVM models: one with the full training data and another with only moral-labeled training data.
For word embedding methods, we use FrameAxis Python package (Kwak et al., 2021) to compute anchor micro-frames based on different MFDs with the word2vec embedding model (Mikolov et al., 2013). For each MFD, we generate the corresponding micro moral frames from the vocabulary in its class. We then compute and aggregate word contributions to each microframe in the document, and label the document’s moral class by identifying significant microframes through comparison with a null model.
For language models, we apply pre-trained language models MoralBert [^5] Accessed from https://huggingface.co/vjosap and Mformers [^6] Accessed from https://huggingface.co/joshnguyen from HuggingFace with their default settings and no additional fine-tuning.
Local Language Lexicons
We apply C-MFD2 (Cheng and Zhang, 2023) to evaluate the performance of the local language dictionary approach. We test two techniques for locally-developed MF lexicons: word counts and embeddings. For the word embedding method, we test two approaches with the fastText model (Grave et al., 2018). One involves measure simple semantic similarity. Words in C-MFD2 are grouped into five pseudo-documents based on their MF labels, each serving as anchor frames. MF values are then determined by calculating the semantic distance between the text to be classified and these anchor frames. The other uses FrameAxis to construct anchor frames. As C-MFD2 does not have the virtue/vice dimension which is essential to calculate micro-frames in FrameAxis, we automatically assign this dimension using the RoBERTa-based Chinese sentiment model c2-roberta-base-finetuned-dianping-chinese from HuggingFace [^7] Accessed from https://huggingface.co/liam168/c2-roberta-base-finetuned-dianping-chinese.
Multilingual Language Models
We select XLM-T  [^8] Accessed from https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base as the base model to test transfer learning on the multilingual language model approach (Barbieri, Anke, and Camacho-Collados, 2022). Fine-tuned on 198 million multilingual tweets on the XLM-RoBERTa architecture—originally trained on 2.5 TB of Common Crawl data (Conneau et al., 2019)—this model is particularly well-suited for analyzing online content. Also, previous research indicates that it has strong performance in cross-language deductive coding tasks such as hate speech detection, compared to other multilingual language models (Röttger et al., 2022).
[…]
After fine-tuning with English annotated data, we further fine-tune each base model with the 80% of CCV dataset in order to test the data-efficiency as in Röttger et al. (2022). We incrementally train models with batches of additional CCV data, with each batch containing 100 annotated records.
Large Language Models
For the LLMs approach, we select the Llama3.1-8b instruct model, an open-source multilingual LLM developed by Meta with a reasonable balance between model performance and computational cost. Compared to closed-source models like GPT-4, Llama3.1 offers greater control, flexibility, transparency, and reproducibility—all of which are important in human-value measurement tasks.
We first test LLMs with prompt-engineering and few-shot learning. Next, we apply the same fine-tuning process used for XLM-T to Llama3.1-8b. Notably, an additional round of data augmentation is performed afterwards, where all English annotations are machine-translated into Chinese using the Google Translate API. Fine-tuning is conducted with the unsloth package using 4-bit quantization on a single NVIDIA L40S GPU. Additionally, we test data efficiency with 20 batches of Chinese annotated items from the CCV dataset. A conservative under-sampling strategy is applied as well with each batch containing 50 records evenly distributed across the five classes.
Results
Machine Translation
As shown in Table 2, the performance of machine translation generally fall short in evaluation. With the MFV benchmarking dataset, the lexicon method MFD2 shows the best performance with a weighted F1 score of $0.60$. In the reverse-annotated CCS dataset, a simple SVM model with lexicon features outperforms other measurements ($F1=0.74$), but the model coverage is relatively low at only 23%. In the real-word CCV dataset, the deep learning model Mformer exhibits the strongest performance ($F1=0.47$) and maintains comparable results across the other two benchmark datasets. Note that although some MF classes in Mformer displayed good performance (i.e., care/harm, $F1=0.72$), some fine-grained MF measurements are very poor. For example, Mformer’s prediction on “loyalty” ($F1=0.21$) is worse than random guessing baseline ($F1=0.23$) in the cross-language evaluation setting.
Auth Care Fair Loya Sanc Acc Cov Fw Fm
MFV
Baseline 0.28 0.30 0.13 0.18 0.11 0.23 1.00 0.23 0.20
MFD 0.77 0.29 0.00 0.55 0.00 0.56 0.28 0.55 0.32
MFD2 0.69 0.50 0.75 0.59 0.33 0.59 0.46 0.60 0.57
eMFD 0.39 0.58 0.31 0.22 0.44 0.42 1.00 0.41 0.39
MS 0.00 0.00 0.00 0.33 0.31 0.17 0.20 0.13 0.13
FA+MFD 0.39 0.06 0.00 0.29 0.31 0.27 1.00 0.21 0.21
FA+MFD2 0.19 0.18 0.06 0.43 0.17 0.22 1.00 0.21 0.21
FA+eMFD 0.07 0.47 0.19 0.40 0.17 0.31 1.00 0.28 0.26
svm+MS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
svm+MS* 0.07 0.32 0.21 0.09 0.00 0.19 1.00 0.16 0.14
MoralBert 0.27 0.55 0.50 0.10 0.30 0.38 1.00 0.36 0.34
MFormer 0.65 0.65 0.72 0.31 0.40 0.58 1.00 0.57 0.55
CCS
Baseline 0.23 0.27 0.18 0.17 0.16 0.21 1.00 0.21 0.22
MFD 0.47 0.62 0.67 0.18 0.84 0.54 0.6 0.54 0.55
MFD2 0.55 0.71 0.71 0.61 0.71 0.66 0.75 0.66 0.66
eMFD 0.48 0.52 0.48 0.44 0.25 0.47 0.98 0.45 0.43
MS 0.16 0.11 0.16 0.18 0.21 0.16 0.17 0.16 0.16
FA+MFD 0.45 0.38 0.31 0.26 0.47 0.38 1.00 0.38 0.37
FA+MFD2 0.39 0.57 0.57 0.39 0.43 0.47 1.00 0.47 0.47
FA+eMFD 0.33 0.42 0.28 0.17 0.22 0.30 1.00 0.30 0.28
svm+MS 0.68 0.71 0.77 0.85 0.61 0.74 0.23 0.74 0.72
svm+MS* 0.21 0.39 0.28 0.15 0.14 0.26 1.00 0.25 0.23
MoralBert 0.41 0.63 0.58 0.64 0.49 0.57 1.00 0.55 0.55
MFormer 0.65 0.71 0.70 0.71 0.71 0.69 1.00 0.69 0.70
CCV
Baseline 0.18 0.40 0.16 0.23 0.03 0.27 1.00 0.27 0.20
MFD 0.33 0.40 0.43 0.3 0.00 0.34 0.47 0.35 0.29
MFD2 0.23 0.62 0.43 0.27 0.06 0.43 0.72 0.43 0.32
eMFD 0.11 0.59 0.36 0.20 0.02 0.40 1.00 0.36 0.26
MS 0.17 0.21 0.23 0.23 0.09 0.20 0.36 0.21 0.19
FA+MFD 0.29 0.30 0.27 0.31 0.06 0.28 1.00 0.29 0.25
FA+MFD2 0.07 0.49 0.35 0.30 0.11 0.34 1.00 0.34 0.27
FA+eMFD 0.04 0.56 0.31 0.24 0.08 0.38 1.00 0.34 0.25
svm+MS 0.28 0.59 0.49 0.27 0.18 0.47 0.17 0.44 0.36
svm+MS* 0.12 0.35 0.25 0.12 0.03 0.23 1.00 0.23 0.17
MoralBert 0.15 0.62 0.47 0.24 0.10 0.45 1.00 0.41 0.32
MFormer 0.23 0.72 0.52 0.21 0.14 0.50 1.00 0.47 0.36
Table 2: Established English moral foundation measurements applied to machine translated text. Acc, Cov, Fw and Fm refers to accuracy, coverage, F1 weighted and F1 macro respectively. Baseline, FA and MS in the first column represent random guessing, FrameAxis and MoralStrength. The best performing methods for each dataset are in bold. MoralStrength models trained with no non-moral labeled data are marked with a star (*). The Coverage column measures the percentage of all moral labels in the prediction.
Local Language Lexicons
Local language lexicon approaches demonstrate similarly moderate performance with word-count methods. Table 3 shows that across all benchmarking datasets, C-MFD2 consistently outperforms other methods although its performance is generally still poor. In the real-world CCV dataset, C-MFD2 achieves the best performance ($F1=0.43$), which is comparable to the machine translation approach with Mformer. We find multilingual word embedding model fastText with FrameAxis framework do not improve the cross-language measurement performance in this task.
Auth Care Fair Loya Sanc Acc Cov Fw Fm
MFV
Baseline 0.28 0.30 0.13 0.18 0.11 0.23 1.00 0.23 0.20
cMFD2 0.48 0.18 0.67 0.29 0.22 0.39 0.40 0.42 0.37
FT 0.44 0.00 0.00 0.12 0.18 0.30 1.00 0.16 0.15
FT+FA 0.08 0.00 0.00 0.30 0.00 0.19 1.00 0.08 0.08
CCS
Baseline 0.23 0.27 0.18 0.17 0.16 0.21 1.00 0.21 0.22
cMFD2 0.74 0.76 0.73 0.68 0.74 0.74 0.76 0.73 0.73
FT 0.44 0.32 0.35 0.39 0.37 0.39 0.97 0.37 0.37
FT+FA 0.14 0.05 0.15 0.29 0.02 0.20 0.97 0.12 0.13
CCV
Baseline 0.18 0.40 0.16 0.23 0.03 0.27 1.00 0.27 0.20
cMFD2 0.26 0.63 0.45 0.28 0.09 0.45 0.64 0.43 0.34
FT 0.28 0.17 0.09 0.01 0.03 0.20 0.98 0.13 0.12
FT+FA 0.02 0.03 0.04 0.37 0.00 0.23 0.98 0.10 0.09
Table 3: The performance of C-MFD2 in cross-language MF measurement. FT and FA in the first column represent FastText and FrameAxis. The best performing methods for each dataset are in bold. Other abbreviations are the same as those in Table 2.
Multilingual Language Models
The multilingual language model XLM-T, trained on the same English annotated data, shows a moderate but reduced performance compared to the monolingual English model Mformer. In Table 4, XLM-T shows moderate performance across all benchmarking datasets ($F1_{MFV}=0.71,F1_{CCS}=0.71,F1_{CCV}=0.63$) outperforming both machine translation and local lexicon approaches. In addition, its performance is generally consistent across the five foundation values, showing a better reliability in the fine-grained MF measurements. This consistency is likely due to training five separate classifiers. In the real-world CCV dataset, the fine-tuned XLM-T model has moderate performance in four out of five foundation values with an average F1 score of $0.63$. The “authority” model has lower performance with $F1=0.46$.
Auth Care Fair Loya Sanc XLM-T Avg
MFV
0 0.70 0.81 0.76 0.88 0.85 0.80
1 0.27 0.41 0.15 0.54 0.15 0.30
Acc 0.58 0.71 0.62 0.81 0.74 0.69
Fm 0.49 0.61 0.45 0.71 0.50 0.55
Fw 0.58 0.69 0.68 0.82 0.77 0.71
CCS
0 0.66 0.76 0.79 0.91 0.89 0.80
1 0.19 0.44 0.02 0.54 0.51 0.34
Acc 0.52 0.66 0.65 0.85 0.82 0.70
Fm 0.42 0.60 0.4 0.73 0.70 0.57
Fw 0.55 0.67 0.65 0.85 0.83 0.71
CCV
0 0.51 0.63 0.74 0.78 0.75 0.68
1 0.24 0.58 0.20 0.41 0.10 0.31
Acc 0.40 0.60 0.61 0.68 0.61 0.58
Fm 0.37 0.60 0.47 0.60 0.43 0.49
Fw 0.46 0.61 0.65 0.70 0.73 0.63
Table 4: Multilingual language models for moral foundation measurement. XLM-T Avg refers to the average F1 score across five foundation models.
In the follow-up evaluation of batch training with local language data, we observe improved model performance as the volume of local language training data increases. As shown in Figure 1, the fine-tuned XLM-T models achieve higher F1 scores with more locally labeled data. For example, the F1 score of the “loyalty” model rose from $0.62$ to $0.80$ with 22 batches of local-language data. Similar increasing patterns are observed across all five models.
However, the amount of annotated non-English data required to reach reliable classification thresholds for MFs is much more than for hate speech detection (Röttger et al., 2022). On average, 16 batches are needed to reach the F1 score of $0.70$ and 49 batches to reach $0.80$. The “care” model, which is the best performing model of the five, requires six additional batches to reach an F1 score of $0.70$ and 27 batches to reach $0.80$. There are 10 batches of data available for the “sanctity” model, and it shows no significant improvement even once all 10 batches are used for training.
It is worth noting that limited local language annotated data may also negatively impact the model’s performance. For example, the “fairness” model’s F1 score initially drops from $0.50$ to $0.46$ when fine-tuned with local-language data. It only begins to stabilize and improve after 11 batches. In sum, while moderate performance can be achieved with local-language annotations, considerably more data is required to attain robust classification performance with the multilingual language model approach in cross-language MF measurements.
Figure 1: Accumulated fine-tuning of the multilingual language model XLM-T with local language annotated data from the CCV dataset. Each batch includes 100 items.
Large Language Models
Table 5 shows the results of using Llama3.1-8b. With only prompt engineering, Llama3.1-8b already performs better than the machine translation and local language lexicon approaches across all benchmarking datasets. After fine-tuning with annotated English-language data, it immediately achieves strong performance on the MFV ($F1=0.81$) and CCS ($F1=0.82$) datasets, and moderate performance on the CCV ($F1=0.65$) dataset. This performance exceeds that of XLM-T with the same English-language training data ($F1=0.63$).
The model is fine-tuned with a data augmentation strategy where the English labeled data are machine translated into Chinese and fed into the LLM again. This step significantly improves the model performance and results in the best performance for all three benchmarking dataset (MFV $F1=0.86$; CCS $F1=0.82$; CCV $F1=0.74$). These F1 scores exceed all other approaches tested in this paper.
Additionally, we notice that English prompts generally outperform non-English prompts on Llama3.1-8b, even when analyzing non-English documents. This difference is more pronounced in the base instruct model but becomes less significant with progressive fine-tuning using annotated data. After fine-tuning with English data, the difference in the F1 scores for English and Chinese prompts on the CCV dataset decreases from 0.11 to 0.08, further reducing to 0.03 with data augmentation and 0.02 when fine-tuned with both English and augmented data.
Moreover, a better LLM base model may further improve the model performance. Table 6 shows that under the same condition, Llama3.1-70b outperforms Llama3.1-8b when prompting in local languages. In the CCV dataset, under a few-shot learning setup with Chinese prompting, the larger LLM ($F1=0.60$) is better than both its smaller counterpart ($F1=0.42$) and the English prompting counterpart ($F1=0.51$).
Auth Care Fair Loya Sanc Acc Cov Fw Fm
MFV
Baseline 0.28 0.30 0.13 0.18 0.11 0.23 1.00 0.23 0.20
[…]
en×$\times$×en 0.89 0.87 0.94 0.57 0.00 0.84 0.41 0.81 0.65
zh×$\times$×en 0.82 0.88 0.75 0.67 0.00 0.79 0.32 0.77 0.62
en×$\times$×zh 0.89 0.86 0.90 0.80 0.00 0.86 0.54 0.85 0.69
zh×$\times$×zh 0.91 0.83 0.91 0.75 0.50 0.85 0.58 0.84 0.78
en×$\times$(en+zh) 0.88 0.89 0.95 0.80 0.00 0.87 0.60 0.86 0.70
zh×$\times$(en+zh) 0.93 0.88 0.91 0.77 0.00 0.87 0.60 0.85 0.70
CCS
Baseline 0.23 0.27 0.18 0.17 0.16 0.21 1.00 0.21 0.22
[…]
en×$\times$×en 0.78 0.85 0.88 0.84 0.68 0.82 0.52 0.82 0.80
zh×$\times$×en 0.75 0.84 0.88 0.80 0.69 0.80 0.45 0.80 0.79
en×$\times$×zh 0.73 0.85 0.84 0.85 0.72 0.81 0.56 0.80 0.80
zh×$\times$×zh 0.74 0.82 0.85 0.86 0.65 0.80 0.56 0.80 0.78
en×$\times$(en+zh) 0.76 0.85 0.86 0.87 0.69 0.82 0.55 0.82 0.81
zh×$\times$(en+zh) 0.74 0.84 0.85 0.87 0.70 0.81 0.57 0.81 0.80
CCV
Baseline 0.18 0.40 0.16 0.23 0.03 0.27 1.00 0.27 0.20
[…]
en×$\times$×en 0.28 0.83 0.68 0.55 0.00 0.66 0.62 0.65 0.47
zh×$\times$×en 0.19 0.82 0.59 0.23 0.00 0.59 0.41 0.57 0.37
en×$\times$×zh 0.39 0.85 0.77 0.69 0.36 0.73 0.68 0.72 0.61
zh×$\times$×zh 0.32 0.81 0.73 0.70 0.36 0.71 0.69 0.69 0.59
en×$\times$(en+zh) 0.42 0.85 0.79 0.72 0.43 0.75 0.72 0.74 0.64
zh×$\times$(en+zh) 0.36 0.84 0.75 0.72 0.31 0.73 0.72 0.72 0.60
Table 5: Llama3.1-8b model with few-shot learning prompts for cross-language moral foundation measurements. The first language refers to the language of prompt while the languages after $\times$ refer to the language(s) of fine-tuning datasets. $\varnothing$ denotes the empty set when no fine-tuning is done. For example, zh×$\times$(en+zh) denotes a Chinese prompt on a model fine-tuned using English and Chinese data. The best performing method per dataset is in bold. Column labels are consistent with Table 2.
Auth Care Fair Loya Sanc Acc Cov Fw Fm
MFV
Baseline 0.28 0.30 0.13 0.18 0.11 0.23 1.00 0.23 0.20
8b en 0.41 0.78 0.40 0.35 0.40 0.55 0.61 0.54 0.47
8b zh 0.56 0.40 0.25 0.00 0.00 0.42 0.36 0.34 0.24
70b en 0.67 0.82 0.52 0.48 0.55 0.66 0.83 0.66 0.61
70b zh 0.73 0.67 0.53 0.56 0.60 0.63 0.64 0.64 0.62
CCS
Baseline 0.23 0.27 0.18 0.17 0.16 0.21 1.00 0.21 0.22
8b en 0.58 0.78 0.71 0.64 0.71 0.70 0.93 0.69 0.68
8b zh 0.66 0.80 0.71 0.63 0.72 0.71 0.82 0.71 0.71
70b en 0.54 0.77 0.63 0.68 0.65 0.67 0.96 0.66 0.65
70b zh 0.56 0.83 0.63 0.71 0.70 0.70 0.89 0.69 0.69
CCV
Baseline 0.18 0.40 0.16 0.23 0.03 0.27 1.00 0.27 0.20
8b en 0.29 0.69 0.47 0.50 0.06 0.54 0.80 0.53 0.40
8b zh 0.29 0.70 0.35 0.16 0.13 0.48 0.61 0.42 0.32
70b en 0.18 0.69 0.52 0.41 0.15 0.55 0.86 0.51 0.39
70b zh 0.18 0.76 0.54 0.59 0.18 0.62 0.63 0.60 0.45
Table 6: Llama3.1 8b and 70b models with few-shot prompting for cross-language moral foundation measurement. The best performing method for each dataset is in bold. Column labels are consistent with Table 2.
Discussion and Limitations
We find traditional approaches, such as machine translation and local language lexicons, may not be the most effective solutions for cross-language moral foundation measurements. Instead, fine-tuning multilingual language models and leveraging LLMs through transfer learning emerge as promising alternatives. Notably, LLMs demonstrate strong performance and data efficiency, making them particularly well-suited for addressing the challenges of cross-language MF analysis. In this section, we extend the discussion to the general human value elements measurements in computational social science research.
Above all, the machine translation approach has proven unreliable in cross-language MF measurement. Even with state-of-the-art English tools, machine translation remains suboptimal in estimating MFs from real-world non-English corpora and, in some cases, performs worse than chance. This issue is particularly concerning for measuring culturally significant human values. For example, within MFT, previous studies highlight East–West cultural differences in emphasis on various foundation values (Graham et al., 2011). Participants from Eastern regions (South Asia, East Asia, and Southeast Asia) generally place a stronger emphasis on “loyalty” compared to their Western counterparts (United States, Canada, United Kingdom, and Western Europe) (Graham et al., 2013). In our test, however, Mformer’s “loyalty” classification on translated data ($F1=0.21$) falls below random chance ($F1=0.23$), suggesting substantial loss of this cultural nuance. Our results caution against over reliance on machine translation in cross-language MF applications and raise broader methodology concerns for its reliability in measuring other cross-language human values.
Researchers should also be cautious with local language lexicons—they perform worse than the machine translation for cross-language MF measurement in some cases. Given the inherent limitations of lexicon-based methods and the extensive resources required to develop them, they are inefficient for cross-language MF measurement, particularly at the document level where semantic complexity is higher. However, culturally specific moral lexicons may still provide valuable insights at the vocabulary level for cross-cultural research, though broader applications should be approached cautiously and with extra validations.
Our evaluation with multilingual language models and LLMs also reveals several key implications. Above all, fine-tuning LLMs is more data-efficient than fine-tuning multilingual language models for cross-language MF measurement. In the CCV benchmarking dataset, XLM-T requires on average more than 2,000 local-language annotated records (20 batches) to achieve strong performance (i.e., $F1=0.75$). In contrast, Llama3.1-8b can reach comparable performance with only English data machine-translated to Chinese and thus no local language labeling.
In addition, data augmentation shows strong potential in enhancing LLMs’ performance on culturally distinct MF values. For example, Llama3.1-8b notably improves on “loyalty” when fine-tuned with English data machine-translated to Chinese ($F1=0.69$), compared to its performance with the original English training data ($F1=0.57$). Such data augmentation may serve as a practical and cost-effective way to improve model performance in cross-language measurements, especially in cases where mass-labeling of local-language data is infeasible due to the resource-intensive process of creating high-quality language-specific labeled datasets.
Our findings also suggest that larger or more advanced LLMs such as Llama3.1-70b may perform better for measuring MFs in non-English corpora. That is, updating the base LLM to a more power model with enhanced multilingual capabilities could further improve the effectiveness of the LLM approach for cross-language MF measurement.
Nevertheless, several limitations of using LLMs for cross-language MF measurement should be acknowledged. Firstly, their performance on fine-grained MF values can vary. For example, although the measurement on some culturally sensitive values like “loyalty” is relatively reliable, others like “authority” and “sanctity” still miss significant cultural nuances. Careful prompt crafting and fine-tuning LLMs for binary instead of multi-label classification task may help mitigate this limitation.
Secondly, the model coverage remains a challenge. Our tested four Llama models in Table 6 on average successfully detect 72.50% of the MF values in the CCV dataset, with coverage varying with the prompting language and fine-tuning data. This issue can be partially attributed to the model safety policies as Llama refuses to respond to certain prompts. Some documents in the benchmarking dataset contain morally controversial content that may prevent Llama from generating outputs—a limitation that is less prominent in multilingual language models and other approaches.
Thirdly, fine-tuning LLMs with limited local-language annotated data might backfire. While training with local-language annotated data is commonly used to incorporate cultural specificity and enhance transfer learning, this approach appears less effective for LLMs when data volume is small. We fine-tuned Llama3.1-8b using 20 batches of CCV data with the same strategy used in XLM-T to test data efficiency. Each batch containing 50 MF labels evenly distributed across five classes. As shown in Figure 2, The initial drop in model performance struggles to recover within the available data for care, fairness and sanctity classifiers. Researchers with limited local-language annotated data should be cautious about potential performance degradation when fine-tuning LLMs for cross-language MF measurements.
Figure 2: Accumulated fine-tuning multilingual language model Llama3.1-8b with local language annotated data from the Chinese CV dataset. Each batch includes 50 items evenly distributed across five classes.
In addition, it is important to acknowledge potential limitations in our mapping scheme. The CCV dataset was originally labeled based on Chinese core values, and the mapping process to moral foundation framing may have led to some cultural information loss.
Finally, we would like to highlight the essential role of human validation in measuring cross-language MFs with LLMs. Given the limitations discussed, human validation is strongly recommended to ensure the reliability of LLM-based cross-language MF measurements. LLMs can effectively assist in this process by providing rationales alongside their classification outputs, supporting human evaluators and facilitating a more efficient assessment of cross-language MF measurements.
In summary, our findings suggest LLMs trained on English, combined with prompt-engineering and data augmentation can achieve robust performance with strong reliability and data efficiency in cross-language MF measurements. Local-language annotations should focus on the culturally distinct moral values to further enhance the performance.
Conclusions
This study examined computational approaches for automatically measuring moral foundation values in non-English texts. It uses Chinese as a case study and leverages established English resources for this cross-language deductive coding task.
It first highlighted the limitations of dictionary and machine translation approaches: while local language dictionaries can support lexicon-level analysis, they often lack the depth needed for complex semantic assessments. Notably, advanced English-based tools applied to machine-translated data outperform local lexicon-based methods, which underscores the limitations of lexicon approaches in this task. The study then explores the potential of transfer learning in language models, showing that both multilingual language models and LLMs demonstrate strong performance, with LLMs performing better and being more data-efficient. It is recommended to select approaches based on the availability of local-language annotated data. When sufficient data is available, a smaller multilingual language model generally yields satisfactory results. Otherwise, LLMs fine-tuned and augmented on English resources offer the best performance, though with higher computational demands. Finally, this study emphasizes the necessity of human validation in cross-language MF measurement. The reasoning capability of LLMs can significantly support this process by generating rationales for classification decisions, assisting human evaluators with more efficient review and reliable assessments.
The findings in this paper provide valuable insights for cross-cultural MF research and shed light on future applications of LLM-assisted deductive coding in multilingual tasks.
Appendix
The appendix presents the mapping scheme of Chinese core values to moral foundation values.

### v3

@@
 Since most resources are developed primarily for English, cross-linguistic applications of moral foundation theory remain limited.
-Using Chinese as a case study, this paper evaluates the effectiveness of applying English resources to machine translated text, local language lexicons, multilingual language models, and large language models (LLMs) in measuring MFs in non-English texts.
+Using Chinese as a case study, this paper evaluates the effectiveness of applying English resources to machine translated text, local language lexicons, multilingual encoder-only language models, and decoder-only large language models (LLMs) in measuring MFs in non-English texts.
 The results indicate that machine translation and local lexicon approaches are insufficient for complex moral assessments, frequently resulting in a substantial loss of cultural information.
-In contrast, multilingual models and LLMs demonstrate reliable cross-language performance with transfer learning, with LLMs excelling in terms of data efficiency.
-Importantly, this study also underscores the need for human-in-the-loop validation of automated MF assessment, as the most advanced models may overlook cultural nuances in cross-language measurements.
+In contrast, language models demonstrate reliable cross-language performance with transfer learning, with LLMs excelling in terms of data efficiency.
+Importantly, this study also underscores the need for human-in-the-loop validation of automated MF assessment, as even the most advanced models may overlook cultural nuances and face potential risks in cultural misalignment.
 The findings highlight the potential of LLMs for cross-language MF measurements and other complex multilingual deductive coding tasks.
 Introduction
-Moral intuitions have long fascinated social scientists, as they help explain a wide range of cognitive and behavioral phenomena across individuals and groups (Effron and Helgason, 2022).
-Moral foundation theory (MFT) is among the most prominent psychology frameworks for understanding the origin and development of human morality (Graham et al., 2013).
-Rooted in moral nativism, MFT argues there are five universal moral foundations—care/harm, fairness/cheating, authority/subversion, loyalty/betrayal, and sanctity/degradation—that transcend languages and cultures and underlie people’s moral judgments and decision-making processes (Graham et al., 2013)  [^1] Each MF includes both virtue and vice dimensions.
-This paper uses the virtue label to represent the entire foundation..
-While some scholars propose other foundations (Haidt, 2012; Atari et al., 2023), these five foundations have received the most empirical validation across domains, languages, and cultures (Iurino and Saucier, 2020).
+Moral intuitions have long fascinated social scientists, as they help explain a wide range of cognitive and behavioral phenomena across individuals and groups (Effron and Helgason 2022).
+Moral foundation theory (MFT) is among the most prominent psychology frameworks for understanding the origin and development of human morality (Graham et al.
+2013).
+Rooted in moral nativism, MFT argues there are five universal moral foundations—care/harm, fairness/cheating, authority/subversion, loyalty/betrayal, and sanctity/degradation—that transcend languages and cultures and underlie people’s moral judgments and decision-making processes (Graham et al.
+2013).
+[^1] Each MF includes both virtue and vice dimensions.
+We use the virtue label to represent the foundation value.
+While some scholars propose other foundations (Haidt 2012; Atari et al.
+2023a), these five foundations have received the most empirical validation across domains, languages, and cultures (Iurino and Saucier 2020).
 A growing body of literature employs MFT to investigate online social behaviors.
 For example, MFT provides a framework for understanding rising political polarization.
-Individuals who prioritize loyalty, authority, and sanctity foundations are more likely to endorse conservative views and engage in polarized political discourse, while those affiliated with liberal ideologies tend to value all MFs more evenly (Koleva et al., 2012; Haidt and Graham, 2007).
+Individuals who prioritize loyalty, authority, and sanctity foundations are more likely to endorse conservative views and engage in polarized political discourse, while those affiliated with liberal ideologies tend to value all MFs more evenly (Koleva et al.
+2012; Haidt and Graham 2007).
 MFT also sheds light on online cultural clashes.
 Atran (2007) identified differing valuations of the sanctity foundation as a key factor in many religious and ideological conflicts.
-Beyond polarization, MFT has been applied to a range of social issues, including climate change (Markowitz and Shariff, 2012), vaccine hesitancy (Amin et al., 2017), anti-abortion views (Koleva et al., 2012), nationalism (Kertzer et al., 2014), violent collectivism (Nussio, 2023), and terrorism (Tamborini et al., 2020).
+Beyond polarization, MFT has been applied to a range of social issues, including climate change (Markowitz and Shariff 2012), vaccine hesitancy (Amin et al.
+2017), anti-abortion views (Koleva et al.
+2012), nationalism (Kertzer et al.
+2014), collective violence (Nussio 2023), and terrorism (Tamborini et al.
+2020).
 Given its broad relevance, measuring MF values in online discourse is essential; yet, automated extraction of MF values from large-scale texts remains challenging, particularly for non-English corpora.
 Like other latent human values, MFs are often conveyed through abstract narratives that vary across languages.
-Also, most computational resources for the measurement of MFs are designed for English content (Hoover et al., 2020; Trager et al., 2022).
-This reliance on English hinders cross-cultural comparative research on MFT and limits theoretical advancement from non-English data (Cheng and Zhang, 2023).
-Although MFT is a universal theory intended to apply across languages and cultures, the limited availability of non-English resources severely restricts its research scope and further development of the theoretical framework (Graham et al., 2013).
+Also, most computational resources for the measurement of MFs are designed for English content (Hoover et al.
+2020; Trager et al.
+2022).
+This reliance on English hinders cross-cultural comparative research on MFT and limits theoretical advancement from non-English data (Cheng and Zhang 2023).
+Although MFT is intended to apply across languages and cultures, the limited availability of non-English resources severely restricts its research scope and further development of the theoretical framework (Graham et al.
+2013).
 In this work, we investigate various computational approaches for cross-language measurement of MFs with a particular focus on data-efficiency.
 We use Chinese as an example and find that (1) MF local language lexicons yield suboptimal performance.
-They are worse than machine translation approaches that utilize established English measurements such as Mformer; (2) multilingual language models can achieve moderate success when trained on annotated English data along with some local language labeled data.
-However, this strategy is less data-efficient in measuring MFs than in other deductive coding tasks (e.g., Röttger et al., 2022).
-(3) LLMs outperform other approaches in cross-language MF measurements in both accuracy and data efficiency.
-Simply fine-tuning and augmenting with English annotated data can achieve strong performance in non-English corpora; (4) nevertheless, this performance is in consistent on different MF values, as LLMs may overlook cultural nuances in cross-language measurements, particularly for culturally distinct values.
-[^2] Code is available at GitHub https://github.com/calvinchengyx/cross-lan-mft-measure.
+They are worse than machine translation approaches that utilize established English measurements such as Mformer.
+(2) Multilingual encoder-only models can achieve moderate success when trained on annotated English data along with some local language labeled data.
+However, this strategy is less data-efficient in measuring MFs than in other deductive coding tasks, such as hate speech detection (Röttger et al.
+2022).
+(3) Decoder-only LLMs outperform other approaches in cross-language MF measurements in both accuracy and data efficiency.
+Simply fine-tuning and augmenting with English annotated data can achieve strong performance on non-English corpora; (4) nevertheless, this performance is inconsistent on different MF values, as LLMs may overlook cultural nuances in cross-language measurements, particularly for culturally distinct values  [^2] Code is available on the project’s GitHub https://github.com/calvinchengyx/cross-lan-mft-measure.
 Related Work
-The unique link between word usage and the expressed moral values provides a theoretical foundation for automated MF measurement from online texts (Brady, Crockett, and Van Bavel, 2020; Gantman and Van Bavel, 2014, 2016).
-Scholars have explored different computational approaches, including dictionaries (Graham, Haidt, and Nosek, 2009; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning (Lan and Paraboni, 2022), deep learning language models (Preniqi et al., 2024; Nguyen et al., 2024) and LLMs (Rathje et al., 2024).
+The unique link between word usage and the expressed moral values provides a theoretical foundation for automated MF measurement from online texts (Brady, Crockett, and Van Bavel 2020; Gantman and Van Bavel 2014, 2016).
+Scholars have explored different computational approaches, including dictionaries (Graham, Haidt, and Nosek 2009; Hopp et al.
+2021), word embeddings (Kwak et al.
+2021; Araque, Gatti, and Kalimeri 2020), machine learning (Lan and Paraboni 2022), deep learning language models (Preniqi et al.
+2024; Nguyen et al.
+2024) and LLMs (Rathje et al.
+2024).
 These computational methods demonstrate great advantages on scalability and labor intensity compared to traditional human annotations.
-However, these approaches are primarily developed for English language, and are not directly applicable to non-English texts due to several major concerns: differences in cultural contexts, the lack of annotated datasets, and limitations in domain generalizability.
-To address the cross-language challenges and bridge knowledge gaps in MF measurements, various computational approaches have been proposed, which can be broadly categorized into two paths: machine translation to English and the development of cross-language measurement tools (Zhuang et al., 2020).
+However, these approaches are primarily developed for English, and are not directly applicable to non-English texts due to several major concerns: differences in cultural contexts, the lack of annotated datasets, and limitations in domain generalizability.
+To address the cross-language challenges and bridge knowledge gaps in MF measurements, various computational approaches have been proposed, which can be broadly categorized into two paths: machine translation to English and the development of cross-language measurement tools (Zhuang et al.
+2020).
 English Centric Machine Translation
 Translation is a widely used technique in cross-language MF measurement.
-For instance, the MFT survey has been translated into over 20 languages for cross-lingual studies (Yilmaz et al., 2016; Nilsson and Erlandsson, 2015).
-For large scale text analysis, the development of multilingual neural machine translation has significantly improved translation quality compared to earlier statistical methods, enhancing context understanding, ambiguity resolution, and fluency (Stasimioti et al., 2020).
-This advancement enables a machine-translated approach to cross-language MF measurement by translating target languages into English and applying established English-based methods (Artetxe, Labaka, and Agirre, 2020).
-Previous research has proposed various established methods for automatically measuring MFs in English, including dictionaries (Graham, Haidt, and Nosek, 2009; Frimer et al., 2019; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning and deep learning models (Preniqi et al., 2024; Nguyen et al., 2024) trained on annotated English-language social media data (Hoover et al., 2020; Trager et al., 2022).
+For instance, the MFT survey has been translated into over 20 languages for cross-lingual studies (Yilmaz et al.
+2016; Nilsson and Erlandsson 2015).
+For large scale text analysis, the development of multilingual neural machine translation has significantly improved translation quality compared to earlier statistical methods, enhancing context understanding, ambiguity resolution, and fluency (Stasimioti et al.
+2020).
+This advancement enables a machine-translated approach to cross-language MF measurement by translating target languages into English and applying established English-based methods (Artetxe, Labaka, and Agirre 2020).
+Established methods for automatically measuring MFs in English include dictionaries (Graham, Haidt, and Nosek 2009; Frimer et al.
+2019; Hopp et al.
+2021), word embeddings (Kwak et al.
+2021; Araque, Gatti, and Kalimeri 2020), machine learning and deep learning models (Preniqi et al.
+2024; Nguyen et al.
+2024) trained on annotated English-language social media data (Hoover et al.
+2020; Trager et al.
+2022).
 Moral foundation dictionaries (MFDs)
 Word-count methods with crafted English moral lexicons are common.
-Scholars have developed four English MFDs.
-The original MFD is an expert-crafted dictionary containing a list of 600 words across five foundation values (Graham, Haidt, and Nosek, 2009).
+There are four common English MFDs.
+The original MFD is an expert-crafted dictionary containing a list of 600 words across five foundation values (Graham, Haidt, and Nosek 2009).
 Frimer et al.
-(2019) then expanded this vocabulary to MFD2 with over 2,000 words by automatically querying similar words with word2vec word embeddings (Mikolov et al., 2013).
-Similarly, Araque, Gatti, and Kalimeri (2020) extended the original MFD to a MoralStrength dictionary with approximately 1,000 English lemmas based on WordNet synsets (Princeton, 2010).
-Compared to MFD2, MoralStrength added a round of crowd-sourced ratings on expanded lemmas.
+(2019) then expanded this vocabulary to MFD2 to over 2,000 words by automatically identifying similar words with word2vec word embeddings (Mikolov et al.
+2013).
+Similarly, Araque, Gatti, and Kalimeri (2020) extended the original MFD to a MoralStrength dictionary with approximately 1,000 English lemmas based on WordNet synsets (Princeton 2010).
+Compared to MFD2, MoralStrength added a round of crowd-sourced ratings on the expanded lemmas.
 Hopp et al.
@@
 Semantic similarity methods using embeddings are another approach.
-To address the limitations of word-count methods, such as context insensitivity and vocabulary coverage (Nguyen et al., 2024), scholars have introduced semantic similarity methods.
+To address the limitations of word-count methods, such as context insensitivity and vocabulary coverage (Nguyen et al.
+2024), scholars have introduced semantic similarity methods.
 For example, Kwak et al.
@@
 Target documents are then converted into vectors using word embedding models, and their MF values are determined by comparing to the micro-frames.
-This method has been used to extract MF values from various online texts (e.g., Mokhberian et al., 2020; Jing and Ahn, 2021).
+This method has been used to extract MF values from various online texts (e.g., Mokhberian et al.
+2020; Jing and Ahn 2021).
 Moral language models
 Supervised classification models with annotated English-language training data have also been used.
-With advancements in language models and efforts to create human-labeled MF training datasets (Hoover et al., 2020; Trager et al., 2022), recent studies have demonstrated the potential of fine-tuning language models.
+With advancements in language models and efforts to create human-labeled MF training datasets (Hoover et al.
+2020; Trager et al.
+2022), recent studies have demonstrated the potential of fine-tuning language models.
 For example, Preniqi et al.
 (2024) fine-tuned a BERT-based classifier MoralBert with large-scale annotated English data and achieved state-of-the-art performance.
-To address generalizability limitations in out-of-domain datasets (Liscio et al., 2022), Nguyen et al.
+To address generalizability limitations in out-of-domain datasets (Liscio et al.
+2022), Nguyen et al.
 (2024) proposed another language model Mformer, and reported superior performance compared to other established English methods in evaluations.
 Although machine translation offers several advantages in cross-language MF measurement, including interpretability, scalability, efficiency and accessibility, it also faces significant limitations.
-First, translation quality varies across languages and domains (Ranathunga et al., 2023).
-In some low-resource languages like Tamil, machine translation often makes errors in translating domain terms, polysemous words, and contains repetitions for semantically similar terms (Ramesh et al., 2021).
-Second, it often fails to retain non-propositional information, such as emotion nuances.
-This can lead to emotion loss, toning down, or amplification across languages, introducing bias in subsequent analyses (Troiano, Klinger, and Padó, 2020).
-Third, machine translation struggles to capture cultural elements, which is a major concern for cross-cultural and comparative research (Haidt, 2012).
-It often shows limited performance with rare or culture-specific words, idiomatic phrases, and metaphor recognition (Dorothy, Keiko, and Vladimir, 2019).
+First, translation quality varies across languages and domains (Ranathunga et al.
+2023).
+In some low-resource languages like Tamil, machine translation often makes errors in translating domain terms, polysemous words, and contains repetitions for semantically similar terms (Ramesh et al.
+2021).
+Second, it often fails to retain non-propositional information, such as emotional nuances.
+This can lead to a loss of emotions, toning down, or amplification across languages, introducing bias in subsequent analyses (Troiano, Klinger, and Padó 2020).
+Third, machine translation struggles to capture cultural elements, which is a major concern for cross-cultural and comparative research (Haidt 2012).
+It often shows limited performance with rare or culture-specific words, idiomatic phrases, and metaphor recognition (Dorothy, Keiko, and Vladimir 2019).
 Thus, it remains unclear whether machine translation is a reliable method for measuring cross-language MFs.
@@
 A second path is to develop cross-language tools, where scholars create computational MF resources tailored to local languages.
-Common cross-language measurements include local language dictionaries, task-specific multilingual language models, and LLMs.
+Common cross-language measurements include local language dictionaries, task-specific encoder-only language models, and LLMs.
 Local language dictionaries
-Due to the efficiency at scale and multilingual capabilities, local language dictionaries are widely used to estimate MF values from non-English texts (Hopp et al., 2021).
-Developing local language MFDs generally involves three steps: (1) translating English MFDs to target languages; (2) adding culturally specific and non-translatable vocabulary; and (3) validating with native speakers and local language corpora.
-Several extensive non-English MFDs have been developed and validated in Turkish (Alper et al., 2020), Japanese (Matsuo et al., 2019), Portuguese (Carvalho et al., 2020), and Chinese (Cheng and Zhang, 2023).
+Due to the efficiency at scale and multilingual capabilities, local language dictionaries are widely used to estimate MF values from non-English texts (Hopp et al.
+2021).
+Developing local language dictionaries generally involves three steps: (1) translating English dictionaries to target languages; (2) adding culturally specific and non-translatable vocabulary; and (3) validating with native speakers and local language corpora.
+Several extensive non-English MFDs have been developed and validated in Turkish (Alper et al.
+2020), Japanese (Matsuo et al.
+2019), Portuguese (Carvalho et al.
+2020), and Chinese (Cheng and Zhang 2023).
 Despite the abovementioned advantages, the dictionary approach still faces the inherent limitations of general bag-of-words methods.
@@
 We also test a semantic similarity approach using the FrameAxis architecture and cross-language word embedding models.
-Multilingual Language Models
+Multilingual encoder-only models
 To overcome the limitations of bag-of-words methods, literature has suggested machine learning and deep learning approaches.
-A primary challenge with these methods is the scarcity of annotated data in local languages for model training (e.g., Ji et al., 2024; Nguyen et al., 2024).
+A primary challenge with these methods is the scarcity of annotated data in local languages for model training (e.g., Ji et al.
+2024; Nguyen et al.
+2024).
 Therefore, scholars have adopted transfer learning techniques that leverage English-annotated resources for cross-language classifier development.
-Two major transfer learning strategies are commonly proposed: (1) machine-translating annotated English-language data into local languages to train monolingual models (Schuster et al., 2019), or (2) using annotated English-language data to train multilingual models (Barriere and Balahur, 2020).
-Given the strong performance of multilingual language models in deductive coding tasks such as sentiment analysis (e.g., Barriere and Balahur, 2020) and hate speech detection (e.g., Röttger et al., 2022), this paper focuses on the second transfer learning strategy and evaluates multilingual language models for cross-language MF measurement.
+Two major transfer learning strategies are commonly proposed: (1) machine-translating annotated English-language data into local languages to train monolingual models (Schuster et al.
+2019), or (2) using annotated English-language data to train multilingual encoder-only models (Barriere and Balahur 2020).
+Multilingual models have demonstrated strong performance in deductive coding tasks, such as sentiment analysis (e.g., Barriere and Balahur 2020) and hate speech detection (e.g., Röttger et al.
+2022).
+Nevertheless, they also have shown language bias in morality classification tasks, as pre-trained multilingual encoder-only language models often display distinct moral directions across languages (Hämmerl et al.
+2023).
+This paper focuses on the second transfer learning strategy and evaluates multilingual encoder-only models for cross-language MF measurement.
 Large language models (LLMs)
-The rise of LLMs provides an alternative for cross-language MF measurement.
-LLMs show exceptional zero/few-shot learning capability, enabling them to directly label human values out-of-the-box, which is particularly valuable for tasks with limited human annotated data (Ziems et al., 2024).
-They also have demonstrated strong performance in measuring various human values, including emotions, stance, and political ideology (Ziems et al., 2024; Gilardi, Alizadeh, and Kubli, 2023; Törnberg, 2023).
-Notably, LLMs sometimes are not as good as specialized fine-tuned language models (Amin, Cambria, and Schuller, 2023; Preniqi et al., 2024), which may be due to the lack of explicit, colloquial definitions of the target human values (Ziems et al., 2024).
+The rise of decoder-only LLMs provides an alternative for cross-language MF measurement.
+LLMs show exceptional zero/few-shot learning capability, enabling them to directly label human values out-of-the-box, which is particularly valuable for tasks with limited human annotated data (Ziems et al.
+2024).
+They also have demonstrated strong performance in measuring various human values, including moral reasoning tasks (Ziems et al.
+2024; Agarwal et al.
+2024).
+Notably, LLMs sometimes are not as good as specialized fine-tuned language models (Amin, Cambria, and Schuller 2023; Preniqi et al.
+2024), which may be due to the lack of explicit, colloquial definitions of the target human values (Ziems et al.
+2024).
 MFT’s well-established conceptual framework may help address this limitation.
-Not only are LLMs pre-trained on rich MFT literature (e.g., Abdulhai et al., 2023), but MFT also offers clear guidance for crafting clear and effective prompts.
-Additionally, LLMs trained on vast multilingual data exhibit promising capabilities to handle cross-language measurements (Ahuja et al., 2023).
-Despite the strengths in accessibility, efficiency, multilingualism, and reasoning, there are also some concerns for LLMs’ performance in cross-language MF measurement.
-First, LLMs exhibit a substantial degree of subordinate multilingualism, displaying proficiency in some languages but not others (Zhang et al., 2023), which has a strong correlation with the proportion of those languages in the pre-training corpus (Li et al., 2024).
-Second, there are potential language biases, particularly in human-value relevant coding tasks (Kirk et al., 2024).
-For example, non-English prompts are more likely to generate malicious responses compared to English prompts (Shen et al., 2024).
-LLMs may also have moral bias across languages, as pre-trained multilingual language models often display distinct moral directions across languages (Hämmerl et al., 2023).
-Third, different LLMs exhibit varying baseline moral tendencies (Ji et al., 2024).
-For instance, GPT-3’s MF preferences align more closely with politically conservative individuals when minimal prompt engineering is used in zero-shot learning (Abdulhai et al., 2023).
-These moral tendencies, however, are highly sensitive to prompting, with different prompting strategies significantly influencing classification outcomes in MF measurements (Abdulhai et al., 2023).
-Thus, it is unclear how LLMs performs in cross-language MF measurement tasks. This paper selects a cutting-edge, open-source model—Llama3.1 (Meta, 2024) to evaluate the LLMs approach.
+Not only are LLMs pre-trained on rich MFT literature (e.g., Abdulhai et al.
+2023), but MFT also offers clear guidance for crafting clear and effective prompts.
+Additionally, LLMs trained on vast multilingual data exhibit promising capabilities to handle cross-language measurements (Ahuja et al.
+2023).
+Despite the strengths in accessibility, efficiency, multilingualism, and reasoning, there are also some concerns in using LLMs’ in cross-language MF measurement.
+First, LLMs exhibit a substantial degree of subordinate multilingualism, displaying proficiency in some languages but not others (Zhang et al.
+2023), which has a strong correlation with the proportion of those languages in the pre-training corpus (Li et al.
+2024).
+Second, there are potential language biases, particularly in human-value relevant coding tasks (Kirk et al.
+2024; Yu et al.
+2024).
+For example, non-English prompts are more likely to generate malicious responses compared to English prompts (Shen et al.
+2024).
+Third, different LLMs exhibit varying baseline moral tendencies (Ji et al.
+2024).
+For instance, GPT-3’s MF preferences align more closely with politically conservative individuals when minimal prompt engineering is used in zero-shot learning (Abdulhai et al.
+2023).
+These moral tendencies, however, are sensitive to prompting, with different prompting strategies significantly influencing classification outcomes in MF measurements (Abdulhai et al.
+2023; Agarwal et al.
+2024).
+Thus, it is unclear how LLMs performs in cross-language MF measurement tasks. This paper selects a cutting-edge, open-source model—Llama3.1 (Meta 2024) to evaluate on LLM approach.
 Data
@@
 We used three human annotated datasets—moral foundation vignettes (MFV), Chinese moral scenarios (CCS) and Chinese core values (CCV), to benchmark the performance of different cross-language MF measurement approaches.
-And we used three English annotated MFT datasets—Reddit Corpus, Twitter Corpus and news corpus—for model training and fine-tuning, aligning with the work of Nguyen et al.
-(2024).
+We also used three English annotated MFT datasets for model training and fine-tuning.
 MFV CCS CCV EN Total
@@
 Moral Foundation Vignettes (MFV)
-MFV is a list of social behaviors constructed by psychologists based on MFT, describing the violations of specific moral values from a third-party perspective (Clifford et al., 2015).
-It has been widely validated and used to assess measures of moral judgment (e.g., Atari et al., 2023; Curry, Chesters, and Van Lissa, 2019).
-The original MFV is in English; we manually translate it to Chinese and ensure the cultural context is preserved.
-MFV is used as a universal stimulus to benchmark the performance of different cross-language MF measurement approaches on an expert-crafted sample.
+MFV is a list of social behaviors constructed by psychologists based on MFT, describing the violations of specific moral values from a third-party perspective (Clifford et al.
+2015).
+It has been widely validated and used to assess measures of moral judgment (e.g., Graham, Haidt, and Nosek 2009; Kivikangas et al.
+2021; Ji et al.
+2024).
+We used the MFV as an expert-crafted benchmark stimulus to evaluate the baseline performance of different moral foundation measurement approaches.
+Since the original MFV is in English, we followed a careful translation process to ensure cultural neutrality in Chinese.
+First, a native Chinese speaker translated the vignettes with minor modifications to preserve cultural appropriateness.
+Then, two additional native speakers reviewed the translations to assess whether the scenarios remained representative and meaningful in Chinese cultural contexts.
+This ensured that the translated vignettes could serve as a culturally neutral benchmark for cross-method comparability.
+[^3] The translated vignettes are available on the project’s GitHub page.
 Chinese Moral Scenarios (CCS)
-CCS is list of moral scenarios written by laypeople to describe their intuitive understanding of MF values (Cheng and Zhang, 2023).
-It is written by 202 native Chinese speakers describing situations that either violate or respect specific MF values, with minimal training in MFT literacy.
-This reverse-annotation method is commonly used for validating MF measurements (e.g., Cheng and Zhang, 2023; Frimer et al., 2019; Matsuo et al., 2019).
+CCS is list of moral scenarios written by 202 native Chinese to describe their intuitive understandings of MF values (Cheng and Zhang 2023).
+This reverse-annotation method is commonly used for validating MF measurements (e.g., Cheng and Zhang 2023; Frimer et al.
+2019; Matsuo et al.
+2019).
 It incorporates culturally specific content, reflecting native speakers’ natural and intuitive understanding of MFT in a real-word context.
 Chinese Core Values (CCV)
-CCV is a human-annotated real-world dataset, including 6,994 sentences collected from four local Chinese news websites.
-[^3] CCV includes data from the CMOS corpus (Peng et al., 2021), China Cultural and Ethical Website wenming.cn, Youth Patriotism News agzy.youth.cn, and Sohu News news.sohu.com.
-The dataset is annotated by three native Chinese speakers based on the Chinese core socialist moral value coding scheme, which is highly correlate with the five universal moral foundation values (Liu et al., 2022).
+CCV is a human-annotated real-world dataset, including 6,994 sentences collected from four local news websites.
+[^4] CCV includes data from the CMOS corpus (Peng et al.
+2021), China Cultural and Ethical Website wenming.cn, Youth Patriotism News agzy.youth.cn, and Sohu News news.sohu.com The dataset is annotated by three native Chinese speakers based on the Chinese core socialist moral value coding scheme, which is highly correlate with the five universal moral foundation values (Liu et al.
+2022).
 The original CCV dataset includes eight labels, [^5] The Chinese core socialist moral values include civility, justice, equality, rule of law, patriotism, dedication, integrity, and friendship.
-The curated mapping scheme to the five moral foundation values is in the Appendix which are qualitatively evaluated and re-mapped to the universal five foundation values following an expert-curated mapping scheme.
+Table 3 in the Appendix showed the curated mapping scheme.
+we employed five Chinese native speakers with postgraduate degrees to re-label the action categories in CCV to five moral foundation values, excluding vice and virtue.
+The mapped values were determined by majority vote.
 We sample 20% of the CCV as the primary benchmarking dataset to evaluate the performance across approaches stratifying on the values.
-The remaining 80% is used as local-language annotated data to test the data-efficiency fine-tuning language models.
+The remainder is used as training data to test the data-efficiency fine-tuning language models.
 All benchmarking documents are single-class labeled.
@@
 English annotated data (EN)
-We use three English annotated MF datasets for transfer learning, including a Twitter corpus (Hoover et al., 2020), a Reddit corpus (Trager et al., 2022) and a news corpus (Hopp et al., 2021), which are widely used in training English MF classifiers and show reliable performance (Nguyen et al., 2024; Preniqi et al., 2024).
+We use three English annotated MF datasets for transfer learning, including a Twitter corpus (Hoover et al.
+2020), a Reddit corpus (Trager et al.
+2022) and a news corpus (Hopp et al.
+2021), which are widely used in training English MF classifiers and show reliable performance (Nguyen et al.
+2024; Preniqi et al.
+2024).
 Methods
@@
 First, we machine translate benchmarking datasets except MFV to Chinese using the Google Cloud API—Basic Translation service.
-Then we estimate the MF values from translated documents with established English measurements, including lexicons MFD (Graham, Haidt, and Nosek, 2009), MFD2 (Frimer et al., 2019), eMFD (Hopp et al., 2021) and MoralStrength (Araque, Gatti, and Kalimeri, 2020); word embeddings with FrameAxis (Kwak et al., 2021); and specialized-fine-tuned language models MoralBert (Preniqi et al., 2024) and Mformer (Nguyen et al., 2024).
-For MFD, MFD2, and eMFD, we calculate word frequencies using the eMFDscore Python package (Hopp et al., 2021).
+Then we estimate the MF values from translated documents with established English measurements, including lexicons MFD (Graham, Haidt, and Nosek 2009), MFD2 (Frimer et al.
+2019), eMFD (Hopp et al.
+2021) and MoralStrength (Araque, Gatti, and Kalimeri 2020); word embeddings with FrameAxis (Kwak et al.
+2021); and specialized-fine-tuned language models MoralBert (Preniqi et al.
+2024) and Mformer (Nguyen et al.
+2024).
+For MFD, MFD2, and eMFD, we calculate word frequencies using the eMFDscore Python package (Hopp et al.
+2021).
 For MFD and MFD 1.0, each document’s MF value is determined by the most frequent MF class in the respective dictionary.
@@
 We sum the probabilities and label the document with the class that has the highest sum.
-We use the moralstrength package (Araque, Gatti, and Kalimeri, 2020) for the MoralStrength dictionary.
+We use the moralstrength package (Araque, Gatti, and Kalimeri 2020) for the MoralStrength dictionary.
 We first test its performance of the lexicon features alone with bag-of-word methods; then we train a Support Vector Machine (SVM) model and combine its lexicon features.
 Since English training sets contain many non-moral labels, while the benchmarking dataset contains only moral labels, we train two SVM models: one with the full training data and another with only moral-labeled training data.
-For word embedding methods, we use FrameAxis Python package (Kwak et al., 2021) to compute anchor micro-frames based on different MFDs with the word2vec embedding model (Mikolov et al., 2013).
+For word embedding methods, we use the FrameAxis Python package (Kwak et al.
+2021) to compute anchor micro-frames based on different MFDs with the word2vec embedding model (Mikolov et al.
+2013).
 For each MFD, we generate the corresponding micro moral frames from the vocabulary in its class.
@@
 Local Language Lexicons
-We apply C-MFD2 (Cheng and Zhang, 2023) to evaluate the performance of the local language dictionary approach.
+We apply C-MFD2 (Cheng and Zhang 2023) to evaluate the performance of the local language dictionary approach.
 We test two techniques for locally-developed MF lexicons: word counts and embeddings.
-For the word embedding method, we test two approaches with the fastText model (Grave et al., 2018).
+For the word embedding method, we test two approaches with the fastText model (Grave et al.
+2018).
 One involves measure simple semantic similarity.
@@
 The other uses FrameAxis to construct anchor frames.
-As C-MFD2 does not have the virtue/vice dimension which is essential to calculate micro-frames in FrameAxis, we automatically assign this dimension using the RoBERTa-based Chinese sentiment model c2-roberta-base-finetuned-dianping-chinese from HuggingFace [^7] Accessed from https://huggingface.co/liam168/c2-roberta-base-finetuned-dianping-chinese.
-Multilingual Language Models
-We select XLM-T  [^8] Accessed from https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base as the base model to test transfer learning on the multilingual language model approach (Barbieri, Anke, and Camacho-Collados, 2022).
-Fine-tuned on 198 million multilingual tweets on the XLM-RoBERTa architecture—originally trained on 2.5 TB of Common Crawl data (Conneau et al., 2019)—this model is particularly well-suited for analyzing online content.
-Also, previous research indicates that it has strong performance in cross-language deductive coding tasks such as hate speech detection, compared to other multilingual language models (Röttger et al., 2022).
-[…]
+As C-MFD2 does not have the virtue/vice dimension, which is essential to calculate micro-frames in FrameAxis, we automatically assign this dimension using the RoBERTa-based Chinese sentiment model c2-roberta-base-finetuned-dianping-chinese from HuggingFace.
+[^8] Accessed from https://huggingface.co/liam168/c2-roberta-base-finetuned-dianping-chinese
+Multilingual Encoder-only Models
+We select XLM-T [^9] Accessed from https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base as the base model to test transfer learning on the multilingual encoder-only model approach (Barbieri, Anke, and Camacho-Collados 2022).
+Fine-tuned on 198 million multilingual tweets on the XLM-RoBERTa architecture—originally trained on 2.5 TB of Common Crawl data (Conneau et al.
+2019)—this model is particularly well-suited for analyzing online content.
+Also, previous research indicates that it has strong performance in cross-language deductive coding tasks such as hate speech detection, compared to other encoder-only models (Röttger et al.
+2022).
+We follow Nguyen et al.
+(2024)’s experience on fine-tuning Mformer with some tweaks.
+First, we replace the base architecture from RoBERTa-base (Liu et al.
+2019) with twitter-xlm-roberta-base and tokenization is handled by the model’s built-in tokenizer (Barbieri, Anke, and Camacho-Collados 2022), with token sequences truncated to a maximum of 512.
+Second, we set the learning rate, epochs, and batch size to $52e-$ 5, 3, and 16 respectively following Röttger et al.
+(2022).
+Third, we opt to fine-tune five binary classifiers rather than a single multi-label classifier because binary models generally outperform multi-label models in English MF measurements (Nguyen et al.
+2024).
+[^10] We note that in hate speech detection, Röttger et al.
+(2022) found no significant performance difference between binary and multi-label models when using XLM-T.
+Given that multi-label models require less storage and training resources, they are a viable alternative for future applications.
+Fourth, we adopt a conservative under-sampling strategy in the English annotated training dataset to address class imbalance, establishing a baseline for future improvements.
 After fine-tuning with English annotated data, we further fine-tune each base model with the 80% of CCV dataset in order to test the data-efficiency as in Röttger et al.
@@
 Large Language Models
-For the LLMs approach, we select the Llama3.1-8b instruct model, an open-source multilingual LLM developed by Meta with a reasonable balance between model performance and computational cost.
+For the decoder-only LLMs approach, we select the Llama3.1-8b instruct model, an open-source LLM developed by Meta with a reasonable balance between model performance and computational cost.
 Compared to closed-source models like GPT-4, Llama3.1 offers greater control, flexibility, transparency, and reproducibility—all of which are important in human-value measurement tasks.
@@
 A conservative under-sampling strategy is applied as well with each batch containing 50 records evenly distributed across the five classes.
+Qualitative Analysis
+We conducted qualitative analysis to gain in-depth understanding of the cultural loss in moral foundation measurements across different approaches.
+By randomly sampling 100 mislabeled records, we compared predicted labels to ground truth (i.e., human annotated labels), focusing on MFs that are known with cultural distinctions between English and Chinese (i.e., authority, loyalty, and sanctity).
 Results
@@
 Other abbreviations are the same as those in Table 2.
-Multilingual Language Models
-The multilingual language model XLM-T, trained on the same English annotated data, shows a moderate but reduced performance compared to the monolingual English model Mformer.
+Multilingual Encoder-only Models
+The multilingual encoder-only model XLM-T, trained on the same English annotated data, shows a moderate but reduced performance compared to the monolingual English model Mformer.
 In Table 4, XLM-T shows moderate performance across all benchmarking datasets ($F1_{MFV}=0.71,F1_{CCS}=0.71,F1_{CCV}=0.63$) outperforming both machine translation and local lexicon approaches.
@@
 Fw 0.46 0.61 0.65 0.70 0.73 0.63
-Table 4: Multilingual language models for moral foundation measurement. XLM-T Avg refers to the average F1 score across five foundation models.
+Table 4: Multilingual encoder-only models for moral foundation measurement. XLM-T Avg refers to the average F1 score across five foundation models.
 In the follow-up evaluation of batch training with local language data, we observe improved model performance as the volume of local language training data increases.
@@
 Similar increasing patterns are observed across all five models.
-However, the amount of annotated non-English data required to reach reliable classification thresholds for MFs is much more than for hate speech detection (Röttger et al., 2022).
+However, the amount of annotated non-English data required to reach reliable classification thresholds for MFs is much more than for hate speech detection (Röttger et al.
+2022).
 On average, 16 batches are needed to reach the F1 score of $0.70$ and 49 batches to reach $0.80$.
@@
 It only begins to stabilize and improve after 11 batches.
-In sum, while moderate performance can be achieved with local-language annotations, considerably more data is required to attain robust classification performance with the multilingual language model approach in cross-language MF measurements.
-Figure 1: Accumulated fine-tuning of the multilingual language model XLM-T with local language annotated data from the CCV dataset. Each batch includes 100 items.
-Large Language Models
-Table 5 shows the results of using Llama3.1-8b.
-With only prompt engineering, Llama3.1-8b already performs better than the machine translation and local language lexicon approaches across all benchmarking datasets.
-After fine-tuning with annotated English-language data, it immediately achieves strong performance on the MFV ($F1=0.81$) and CCS ($F1=0.82$) datasets, and moderate performance on the CCV ($F1=0.65$) dataset.
-This performance exceeds that of XLM-T with the same English-language training data ($F1=0.63$).
-The model is fine-tuned with a data augmentation strategy where the English labeled data are machine translated into Chinese and fed into the LLM again.
-This step significantly improves the model performance and results in the best performance for all three benchmarking dataset (MFV $F1=0.86$; CCS $F1=0.82$; CCV $F1=0.74$).
-These F1 scores exceed all other approaches tested in this paper.
-Additionally, we notice that English prompts generally outperform non-English prompts on Llama3.1-8b, even when analyzing non-English documents.
-This difference is more pronounced in the base instruct model but becomes less significant with progressive fine-tuning using annotated data.
-After fine-tuning with English data, the difference in the F1 scores for English and Chinese prompts on the CCV dataset decreases from 0.11 to 0.08, further reducing to 0.03 with data augmentation and 0.02 when fine-tuned with both English and augmented data.
-Moreover, a better LLM base model may further improve the model performance.
-Table 6 shows that under the same condition, Llama3.1-70b outperforms Llama3.1-8b when prompting in local languages.
-In the CCV dataset, under a few-shot learning setup with Chinese prompting, the larger LLM ($F1=0.60$) is better than both its smaller counterpart ($F1=0.42$) and the English prompting counterpart ($F1=0.51$).
+In summary, while moderate performance can be achieved with local-language annotations, considerably more data is required to attain robust classification performance with the multilingual encoder-only model approach in cross-language MF measurements.
+Figure 1: Accumulated fine-tuning of the XLM-T with local language annotated data from the CCV dataset. Each batch includes 100 items.
 Auth Care Fair Loya Sanc Acc Cov Fw Fm
@@
 Column labels are consistent with Table 2.
+Large Language Models
+Table 5 shows the results of the LLMs approach.
+With only prompt engineering, Llama3.1-8b already performs better than the machine translation and local language lexicon approaches across all benchmarking datasets.
+After fine-tuning with annotated English-language data, it immediately achieves strong performance on the MFV ($F1=0.81$) and CCS ($F1=0.82$) datasets, and moderate performance on the CCV ($F1=0.65$) dataset.
+This performance exceeds that of XLM-T with the same English-language training data ($F1=0.63$).
+The model is fine-tuned with a data augmentation strategy where the English labeled data are machine translated into Chinese and fed into the LLM again.
+This step significantly improves the model performance and results in the best performance for all three benchmarking dataset (MFV $F1=0.86$; CCS $F1=0.82$; CCV $F1=0.74$).
+These F1 scores exceed all other approaches tested in this paper.
+Additionally, we notice that English prompts generally outperform non-English prompts on Llama3.1-8b, even when analyzing non-English documents.
+This difference is more pronounced in the base instruct model but becomes less significant with progressive fine-tuning using annotated data.
+After fine-tuning with English data, the difference in the F1 scores for English and Chinese prompts on the CCV dataset decreases from 0.11 to 0.08, further reducing to 0.03 with data augmentation and 0.02 when fine-tuned with both English and translated Chinese data.
+Moreover, a better LLM base model may further improve the model performance.
+Table 6 shows that under the same condition, Llama3.1-70b outperforms Llama3.1-8b when prompting in local languages.
+In the CCV dataset, under a few-shot learning setup with Chinese prompting, the larger LLM ($F1=0.60$) is better than both its smaller counterpart ($F1=0.42$) and the English prompting counterpart ($F1=0.51$).
+Qualitative Analysis of Cultural Nuances
+We selected MFormer and fine-tuned Llama3.1-8b-instruct with English and translated Chinese data as models to analyze cultural loss in translation and non-translation approaches.
+We identified three common types of information loss during the translation process (See Table 4 in the Appendix).
+First, idioms and slang are often mistranslated, as in Example 1, where the Chinese phrase meaning “faked or staged accident for compensation” is simply translated to its superficial meaning.
+Second, contextual meaning.
+As illustrated in Example 2, where the English translation “look towards” misses the cultural connotation of “favoring someone,” resulting in different expressed moral foundations.
+Third, political euphemisms lose their cultural context, as in Example 3, where the English translation fails to convey that the text refers to civil servant behavior.
+These information loss in the translation process often result in mislabeling in MF measurement.
+We further analyzed the non-translational approach to understand how LLMs process cultural nuances.
+Compared to the translational approach, LLMs better recognize slang and idioms (Examples 2 and 13), but demonstrate more subtle cultural misalignment.
+Our analysis revealed several distinct patterns.
+Most notably, LLMs failed to distinguish certain Confucian morality concepts such as filial piety, which encompass family responsibilities to respect and care for elders (Examples 6, 8, 9, and 10).
+Native Chinese speakers tend to labeled these instances as “authority,” whereas language models often only categorized them as “care.” Additionally, we observed misalignment regarding government and local political entities.
+Human annotators tended to label cases involving government (Examples 3, 4) and officials (Example 11) as “loyalty,” while LLMs classified them as “authority” or “fairness.” Finally, examining misalignment in “sanctity” revealed that Chinese annotators frequently associated sanctity with corrupt politicians or role model civil servants (Examples 11 and 12), a distinction rarely captured by LLMs.
 Discussion and Limitations
 We find traditional approaches, such as machine translation and local language lexicons, may not be the most effective solutions for cross-language moral foundation measurements.
-Instead, fine-tuning multilingual language models and leveraging LLMs through transfer learning emerge as promising alternatives.
+Instead, fine-tuning multilingual encoder-only models and leveraging LLMs through transfer learning emerge as promising alternatives.
 Notably, LLMs demonstrate strong performance and data efficiency, making them particularly well-suited for addressing the challenges of cross-language MF analysis.
-In this section, we extend the discussion to the general human value elements measurements in computational social science research.
-Above all, the machine translation approach has proven unreliable in cross-language MF measurement.
-Even with state-of-the-art English tools, machine translation remains suboptimal in estimating MFs from real-world non-English corpora and, in some cases, performs worse than chance.
-This issue is particularly concerning for measuring culturally significant human values.
-For example, within MFT, previous studies highlight East–West cultural differences in emphasis on various foundation values (Graham et al., 2011).
-Participants from Eastern regions (South Asia, East Asia, and Southeast Asia) generally place a stronger emphasis on “loyalty” compared to their Western counterparts (United States, Canada, United Kingdom, and Western Europe) (Graham et al., 2013).
-In our test, however, Mformer’s “loyalty” classification on translated data ($F1=0.21$) falls below random chance ($F1=0.23$), suggesting substantial loss of this cultural nuance.
-Our results caution against over reliance on machine translation in cross-language MF applications and raise broader methodology concerns for its reliability in measuring other cross-language human values.
+Above all, simple machine translation approach has proven suboptimal in cross-language MF measurement.
+This issue is particularly concerning for measuring culturally significant values.
+Machine translation often introduces bias in the rendering of slangs, contextual meanings, and political euphemisms into local languages.
+Moreover, when relying on pretrained English classifiers such as MFormer—which is predominantly trained on English-annotated data—this cultural information loss may be further amplified.
+For instance, MFormer’s classification performance on the “loyalty” foundation in translated data ($F1=0.21$) falls below random chance ($F1=0.23$), indicating a substantial loss of cultural nuance.
+Nevertheless, it is important to note that machine translation models are rapidly evolving.
+Current deficiencies—such as loss of contextual nuance or culturally embedded moral cues—limit their effectiveness in this domain.
+Advances in context-aware and domain-adapted translation models (Jin et al.
+2023; Saunders 2022), particularly those fine-tuned on moral discourse, may mitigate some of these issues, which deserves future exploration.
 Researchers should also be cautious with local language lexicons—they perform worse than the machine translation for cross-language MF measurement in some cases.
 Given the inherent limitations of lexicon-based methods and the extensive resources required to develop them, they are inefficient for cross-language MF measurement, particularly at the document level where semantic complexity is higher.
-However, culturally specific moral lexicons may still provide valuable insights at the vocabulary level for cross-cultural research, though broader applications should be approached cautiously and with extra validations.
-Our evaluation with multilingual language models and LLMs also reveals several key implications.
-Above all, fine-tuning LLMs is more data-efficient than fine-tuning multilingual language models for cross-language MF measurement.
-In the CCV benchmarking dataset, XLM-T requires on average more than 2,000 local-language annotated records (20 batches) to achieve strong performance (i.e., $F1=0.75$).
+However, culturally specific moral lexicons may still provide valuable insights at the vocabulary level for cross-cultural research, though broader applications should be approached with careful validations.
+Fine-tuning LLMs is more data-efficient than fine-tuning multilingual encoder-only models for cross-language MF measurement.
+In the CCV benchmarking dataset, XLM-T requires on average more than 2,000 local-language annotated records (20 batches) to achieve strong performance ($F1=0.75$).
 In contrast, Llama3.1-8b can reach comparable performance with only English data machine-translated to Chinese and thus no local language labeling.
-In addition, data augmentation shows strong potential in enhancing LLMs’ performance on culturally distinct MF values.
-For example, Llama3.1-8b notably improves on “loyalty” when fine-tuned with English data machine-translated to Chinese ($F1=0.69$), compared to its performance with the original English training data ($F1=0.57$).
-Such data augmentation may serve as a practical and cost-effective way to improve model performance in cross-language measurements, especially in cases where mass-labeling of local-language data is infeasible due to the resource-intensive process of creating high-quality language-specific labeled datasets.
-Our findings also suggest that larger or more advanced LLMs such as Llama3.1-70b may perform better for measuring MFs in non-English corpora.
-That is, updating the base LLM to a more power model with enhanced multilingual capabilities could further improve the effectiveness of the LLM approach for cross-language MF measurement.
+Our findings also suggest that larger LLMs such as Llama3.1-70b may perform better for measuring MFs in non-English corpora.
+That is, updating the base LLM to a more power model with enhanced multilingual capabilities could further improve the performance of the LLM approach.
 Nevertheless, several limitations of using LLMs for cross-language MF measurement should be acknowledged.
-Firstly, their performance on fine-grained MF values can vary.
-For example, although the measurement on some culturally sensitive values like “loyalty” is relatively reliable, others like “authority” and “sanctity” still miss significant cultural nuances.
-Careful prompt crafting and fine-tuning LLMs for binary instead of multi-label classification task may help mitigate this limitation.
-Secondly, the model coverage remains a challenge.
-Our tested four Llama models in Table 6 on average successfully detect 72.50% of the MF values in the CCV dataset, with coverage varying with the prompting language and fine-tuning data.
-This issue can be partially attributed to the model safety policies as Llama refuses to respond to certain prompts.
-Some documents in the benchmarking dataset contain morally controversial content that may prevent Llama from generating outputs—a limitation that is less prominent in multilingual language models and other approaches.
-Thirdly, fine-tuning LLMs with limited local-language annotated data might backfire.
-While training with local-language annotated data is commonly used to incorporate cultural specificity and enhance transfer learning, this approach appears less effective for LLMs when data volume is small.
-We fine-tuned Llama3.1-8b using 20 batches of CCV data with the same strategy used in XLM-T to test data efficiency.
+Firstly, the performance on fine-grained MF values can vary.
+Although the measurement on some culturally sensitive values like “loyalty” is relatively reliable, others like “authority” and “sanctity” still miss significant cultural nuances.
+Second, fine-tuning LLMs might backfire.
+Fine-tuning with limited local-language data may degrades model performance.
+While training with local language annotated data is commonly used to incorporate cultural specificity and enhance transfer learning, this approach appears less effective for LLMs in MF measurements task when data volume is small.
+We fine-tuned Llama3.1-8b using 20 batches of CCV data with the same strategy used in XLM-T.
 Each batch containing 50 MF labels evenly distributed across five classes.
-As shown in Figure 2, The initial drop in model performance struggles to recover within the available data for care, fairness and sanctity classifiers.
-Researchers with limited local-language annotated data should be cautious about potential performance degradation when fine-tuning LLMs for cross-language MF measurements.
-Figure 2: Accumulated fine-tuning multilingual language model Llama3.1-8b with local language annotated data from the Chinese CV dataset.
-Each batch includes 50 items evenly distributed across five classes.
-In addition, it is important to acknowledge potential limitations in our mapping scheme.
-The CCV dataset was originally labeled based on Chinese core values, and the mapping process to moral foundation framing may have led to some cultural information loss.
-Finally, we would like to highlight the essential role of human validation in measuring cross-language MFs with LLMs.
+As shown in Figure 2, the initial drop in model performance struggles to recover within the available data for ”care”, ”fairness”, and ”sanctity” values.
+In addition, fine-tuning with English-annotated data may introduce extra cultural bias to LLMs.
+We replicated the same finetuning strategy to an Italian dataset “moralConvITA” (Stranisci et al.
+2021), and observed similar significant performance drops in culturally-nuanced values like ”loyalty” and ”sanctity” (see Appendix).
+This signals the risk of cultural bias in English-centric LLMs.
+Finetuning such LLMs with English annotation data may exacerbate their bias against non-English languages, resulting in more cultural information loss in the measurement task, which echoes concerns about LLMs being shaped by values associated with WEIRD populations (Atari et al.
+2023b).
+These findings highlight the need for culturally-sensitive fine-tuning approaches that preserve cross-cultural moral nuances.
+Figure 2: Accumulated fine-tuning Llama3.1-8b with local language annotated data from the Chinese CV dataset. Each batch includes 50 items evenly distributed across five classes.
+Third, we note that our experiments were conducted in a single non-English language, which may limit the generalizability of our findings.
+Nevertheless, we hope this work can serve as a foundation for future research that extends to a broader range of non-English languages on MF measurement, particularly low-resource ones, where such methodological advancements would be especially valuable.
+Finally, we highlight the essential role of human validation in measuring cross-language MFs with LLMs.
 Given the limitations discussed, human validation is strongly recommended to ensure the reliability of LLM-based cross-language MF measurements.
-LLMs can effectively assist in this process by providing rationales alongside their classification outputs, supporting human evaluators and facilitating a more efficient assessment of cross-language MF measurements.
-In summary, our findings suggest LLMs trained on English, combined with prompt-engineering and data augmentation can achieve robust performance with strong reliability and data efficiency in cross-language MF measurements.
-Local-language annotations should focus on the culturally distinct moral values to further enhance the performance.
+LLMs can effectively assist in this process by providing rationales alongside their classification outputs, supporting human evaluators and facilitating a more efficient result assessment.
 Conclusions
-This study examined computational approaches for automatically measuring moral foundation values in non-English texts.
+This study examined four computational approaches—machine translation, dictionary, multilingual encoder-only models and decoder-only LLMs—for automatically measuring moral foundation values in non-English texts.
 It uses Chinese as a case study and leverages established English resources for this cross-language deductive coding task.
-It first highlighted the limitations of dictionary and machine translation approaches: while local language dictionaries can support lexicon-level analysis, they often lack the depth needed for complex semantic assessments.
+It first highlights the limitations of dictionary and machine translation approaches: while local language dictionaries can support lexicon-level analysis, they often lack the depth needed for complex semantic assessments.
 Notably, advanced English-based tools applied to machine-translated data outperform local lexicon-based methods, which underscores the limitations of lexicon approaches in this task.
-The study then explores the potential of transfer learning in language models, showing that both multilingual language models and LLMs demonstrate strong performance, with LLMs performing better and being more data-efficient.
+The study then explores the potential of transfer learning in language models, showing that both multilingual encoder-only models and LLMs demonstrate strong performance, with LLMs performing better and being more data-efficient.
 It is recommended to select approaches based on the availability of local-language annotated data.
 When sufficient data is available, a smaller multilingual language model generally yields satisfactory results.
-Otherwise, LLMs fine-tuned and augmented on English resources offer the best performance, though with higher computational demands.
-Finally, this study emphasizes the necessity of human validation in cross-language MF measurement.
-The reasoning capability of LLMs can significantly support this process by generating rationales for classification decisions, assisting human evaluators with more efficient review and reliable assessments.
-The findings in this paper provide valuable insights for cross-cultural MF research and shed light on future applications of LLM-assisted deductive coding in multilingual tasks.
+Otherwise, LLMs can serve as a reliable tool.
+We recommend the following steps for applying LLMs to cross-language MF measurement: (1) start with multilingual LLMs and use culturally specific prompt engineering; (2) adopt a binary classification approach for each moral foundation; (3) fine-tune models using available English annotations alongside carefully translated local-language data, while curating out English-centric value cases; and (4) incorporate human validation, especially for culturally distinctive values.
+Findings in this paper provide valuable insights for cross-cultural MF research and shed light on future applications of LLM-assisted deductive coding in multilingual tasks.
 Appendix
-The appendix presents the mapping scheme of Chinese core values to moral foundation values.
+Methods
+We provide additional details on model training and fine-tuning beyond those reported in the method section.
+Python scripts for replication are available on GitHub.
+All models were trained and fine-tuned on identical English annotation datasets: moral foundation Reddit corpus, moral foundation Twitter corpus, and extended moral foundation dictionary news corpus.
+[^11] MFRC https://paperswithcode.com/dataset/mfrc, MFTC https://osf.io/k5n7y/, eMFD https://osf.io/preprints/psyarxiv/924gq˙v1 Computations were performed on a single NVIDIA L40S GPU.
+XLM-T
+Task objective: Binary classification of five moral foundation values.
+Base model: cardiffnlp/twitter-xlm-roberta-base from Hugging Face.
+Hyperparameters: Learning rate: $52e-$ 5.
+Epochs: 3.
+Batch size: 16 (train & eval).
+Weight decay: 0.01.
+Warmup steps: 100.
+Llama3.1
+Task objective: Multiclass classification of five moral foundation values.
+Base model: llama3.1-8b-instruct from Hugging Face.
+Framework: Unsloth with PEFT via LoRA.
+Precision: LoRA applied with 4-bit quantization.
+We used the llama-3 chat template from Unsloth.
+Hyperparameters: Learning rate: $52e-$ 5.
+Epochs: 3.
+Batch size: 128 (train & eval).
+Weight decay: 0.01.
+Max Seq Length: 1024.
+LoRA configurations used default settings: r-16, targeting q, k, v, o, gate, up, and down-proj parameters.
+Random state was set to 3047.
+The English system prompt is in the Appendix.
+Chinese and Italian prompts were translated accordingly and shared on the GitHub.
+Few-shot examples ($N=3$) were purposefully sampled from the benchmarking dataset in the local language.
+Results
+Qualitative Analysis
+Table 4 shows examples of mis-annotated texts in the CCV dataset.
+English translations were generated via Google Translate using the deep-translator Python package.
+MFormer shows results from machine-translated English text, while Llama shows results from original Chinese texts.
+We used the Llama3.1-8b-instruct model fine-tuned on English annotations and their Chinese translations.
+LLMs Approach on Italian
+We tested a similar LLM-based strategy on an Italian dataset “moralConvITA” (Stranisci et al.
+2021), with results shown in Table 7.
+This benchmark comprises 1,724 unique Twitter posts discussing the immigration issue in Italy, labeled by native Italian speakers ($N=8$) according to five moral foundations.
+To minimize annotation ambiguity, we sampled approximately 100 texts with distinct labels for each moral foundation.
+[…]
+We observed similar results as shown in CCV with some noticeable distinctions.
+Finetuning with translated local language annotations shows strong potential in enhancing LLMs’ performance on culturally distinct MF values in CCV.
+For example, Llama3.1-8b notably improves on “loyalty” when fine-tuned with English data machine-translated to Chinese ($F1=0.69$), compared to its performance with the original English training data ($F1=0.57$).
+Such strategy may serve as a practical and cost-effective way to improve model performance in cross-language measurements, especially in cases where mass-labeling of local-language data is infeasible due to the resource-intensive process of creating high-quality language-specific labeled datasets.
+This strategy, however, showed moderate performance on the Italian dataset.
+This difference may be attributed to the smaller Italian benchmark sample size ($N=492$ vs.
+approximately 1,500 in CCV), the unstructured nature of Twitter data (typically brief with limited context), and differences in domain and language complexity.
+The brevity and informality of Italian tweets likely posed greater challenges for accurate moral foundation inference.
+Auth Care Fair Loya Sanc Acc Cov Fw Fm
+baseline 0.20 0.20 0.20 0.20 0.20 0.20 1.00 0.20 0.20
+[…]
+[…]
+en $\times$ en 0.57 0.48 0.43 0.25 0.24 0.44 1.00 0.39 0.39
+it $\times$ en 0.69 0.27 0.09 0.42 0.4 0.42 0.94 0.37 0.37
+en $\times$ it 0.22 0.49 0.32 0.38 0.16 0.34 0.76 0.31 0.31
+it $\times$ it 0.72 0.36 0.09 0.51 0.35 0.46 0.82 0.41 0.41
+en $\times$(en+it) 0.48 0.47 0.37 0.04 0.02 0.37 1.00 0.27 0.27
+it $\times$(en+it) 0.43 0.38 0.16 0.64 0.06 0.39 0.94 0.34 0.33
+Table 7: Llama3.1-8b model for moral foundation measurements on sampled MoralCovnITA dataset.
+The first language refers to the language of prompt while the languages after $\times$×refer to the language(s) of fine-tuning datasets.
+$\varnothing$ denotes no fine-tuning.
+The best performing method is in bold.
+Column labels are consistent with Table 2.
+Then we further qualitative assessed the performance on some moral foundation categories and revealed the cultural misalignment limitation in the finetuning process.
+Fine-tuning with English annotations sometimes produced negative effects.
+Performance on culturally sensitive foundations such as “loyalty” and “sanctity” declined significantly after fine-tuning with English and translated Italian data.
+This signals a concerning cultural misalignment in current LLMs — fine-tuning with English-centric moral foundations may exacerbate inherent cultural biases.
+Our qualitative reading of some mislabeled Italian texts supports this finding.
+For example, Italian understanding of sanctity, influenced by Catholic values and historical context, differs significantly from American conceptions, resulting in poor performance after fine-tuning.
+Future research should further investigate fine-tuning techniques that mitigate cultural misalignment.
+Moreover, we acknowledge the challenges involved in human annotation of moral foundations.
+Both CCV and MoralConvITA datasets have limitations, with annotator disagreements being common — even the same annotator may label identical text differently at different times.
+Although training can reduce inconsistency, maintaining quality control remains difficult at scale.
+Both datasets relied on expert coders, which may introduce labeling biases.
+Additionally, MoralConvITA is subject to topic and text form biases, as it focuses on immigration during a specific period and consists of Twitter posts.
+In contrast, the CCV dataset covers a broader range of topics and primarily includes posts from news articles.
+Consequently, a general performance drop on the Italian dataset is expected.
+These challenges underscore the importance of creating high-quality benchmarking datasets for moral foundation measurement.
+Future research should consider employing demographically representative crowd workers to better capture the distribution of cultural values in local contexts (Hopp and Weber 2021).
+In conclusion, LLMs offer a promising approach for measuring moral foundations in non-English texts when limited local language annotations are available.
+With careful prompting and few-shot learning, LLMs deliver moderate performance compared to human annotation.
+It should be noted that fine-tuning with English annotations though may improve the performance on some moral foundation values, it can introduce additional biases to cultural-sensitive ones, therefore diminish the model performance.
+Figure 3: Mapping Scheme of Chinese Core Value Dataset
+Figure 4: Qualitative analysis of cultural nuances.
+Chinese column refers to the target Chinese text; English Translation column is the result of machine translation using Google Translate.
+Human, MFormer and Llama refer to the labels annotated by human annotators, MFormer, and Llama3.1-8b-instruct model.
+Prompt - English for CCV you are a native Chinese speaker and social science annotator, your task is to label the moral foundation values expressed in the given Chinese documents.
+moral foundation values are the core values that underlie moral reasoning from the moral foundation theory.
+the five moral foundations are: care, fairness, loyalty, authority, and sanctity.
+And they refer to the following moral intuitions, each includes both vice/virtue pairs: - care: related to our long evolution as mammals with attachment systems and an ability to feel (and dislike) the pain of others.
+It underlies the virtues of kindness, gentleness, and nurturance.
+- fairness: related to the evolutionary process of reciprocal altruism.
+It underlies the virtues of justice and rights.
+- loyalty:  our long history as tribal creatures able to form shifting coalitions.
+It is active anytime people feel that its "one for all and all for one." It underlies the virtues of patriotism and self-sacrifice for the group.
+- authority: This foundation was shaped by our long primate history of hierarchical social interactions.
+It underlies virtues of leadership and followership, including deference to prestigious authority figures and respect for traditions.
+- sanctity: This foundation was shaped by the psychology of disgust and contamination.
+It underlies notions of striving to live in an elevated, less carnal, more noble, and more natural way (often present in religious narratives).
+This foundation underlies the widespread idea that the body is a temple that can be desecrated by immoral activities and contaminants (an idea not unique to religious traditions).
+It underlies the virtues of self-discipline, self-improvement, naturalness, and spirituality.
+you should follow the given principles to label the moral foundation values in the give documents: 1.
+identify the moral foundation value only from the 5 given ones.
+2.
+if the document expresses more than 1 foundation value, label all prominent values, but in total should be equal or less than  3 values.
+3.
+provide a brief rationale for the each labelling, which should be less than 20 words.
+4.
+labels the value in english, 5.
+rationales should be in the same lanaguage as the document 6.
+if the document does not express any of the 5 values, label it as ’none’ and provide a brief rationale.
+7.
+if the document can not be labelled into any of the 5 values, label it as ’unknown’ and provide a brief rationale.
+8.
+consider the Chinese cultural context of the document when labelling the values.
+You MUST respond with a brief rationale within 15 words, and the labels.
+save in the dictionary format: {"rationale": "reasons to explain your decision", "labels": "you labels here"} Here are the given documents for your task: